### v2

# Texture-aware and Shape-guided Transformer for Sequential DeepFake Detection

Abstract
Sequential DeepFake detection is an emerging task that aims to predict the manipulation sequence in order. Existing methods typically formulate it as an image-to-sequence problem, employing conventional Transformer architectures for detection. However, these methods lack dedicated design and consequently result in limited performance. In this paper, we propose a novel Texture-aware and Shape-guided Transformer to enhance detection performance. Our method features four major improvements. Firstly, we describe a texture-aware branch that effectively captures subtle manipulation traces with the Diversiform Pixel Difference Attention module. Then we introduce a Bidirectional Interaction Cross-attention module that seeks deep correlations among spatial and sequential features, enabling effective modeling of complex manipulation traces. To further enhance the cross-attention, we describe a Shape-guided Gaussian mapping strategy, providing initial priors of the manipulation shape. Finally, observing that the latter manipulation in a sequence may influence traces left in the earlier one, we intriguingly invert the prediction order from forward to backward, leading to notable gains as expected. Extensive experimental results demonstrate that our method outperforms others by a large margin, highlighting the superiority of our method.
1 Introduction
Recent advancements in deep generative techniques have significantly improved the visual quality of generated faces. One such technique, DeepFake, allows for effortless manipulation of faces, with results that are often imperceptible to the naked eye. This technique has found applications in virtual reality and movie special effects, but it also raises severe security concerns He et al. (2021). To address these concerns, various forensic methods have been proposed to expose DeepFakes, garnering considerable attention.
Conventional forensics methods mainly focus on determining whether a given face image is authentic or forged, which typically falls into the binary classification task. However, with the rapid evolution of face editing tools, the manipulation process has become more versatile and user-friendly, enabling the use of multiple specific operations iteratively until the desired effects are reached. Under this circumstance, existing methods have limitations as they can only identify the final step of manipulation, providing limited insights into the manipulation details and hindering traceability and interpretability. To overcome this issue, a new forensics task called Sequential DeepFake Detection Shao et al. (2022) has emerged recently, aiming to uncover the sequences of face manipulations (see Fig. 1(a,b)). This task expands the scope of DeepFake detection and poses a new challenge to this field.
Figure 1: (a) and (b) describe the difference between conventional DeepFake detection and sequential DeepFake detection. (c) illustrates the process of classic sequential DeepFake detection methods, while (d) represents our proposed method.
Typically, Transformer-based architectures are employed in sequential DeepFake detection to capture traces of different manipulations and predict the manipulation sequences, following the pipeline of image-to-sequence Vinyals et al. (2015). These methods commonly employ a plain Transformer encoder to extract spatial clues and utilize a Transformer decoder to capture the sequential clues (see Fig. 1(c)). Despite these methods showing a certain effectiveness, many critical perspectives are lacking in further exploration.
  1. $\clubsuit$
    Fine-grained Traces Extraction: Generative models often exhibit imperfect semantic disentanglement. One specific manipulation may affect the entire facial structure and leave subtle traces on unintended manipulated regions. To expose the sequential manipulation, existing methods usually utilize a general Transformer encoder to capture those traces. However, this encoder concentrates on the global relationship of the entire faces and may fail to capture the subtle manipulation traces.
  2. $\clubsuit$
    Manipulation Shape Guidance: In the training stage, existing methods enhance the modeling of sequential relation in the decoder by estimating the scale and position of manipulation regions as priors. However, such priors are impractical as the manipulation regions are likely irregular, which could not be well described by only using the scale and position.
  3. $\clubsuit$
    Prediction Order: Existing methods predict the manipulation sequences in a forward order, which is the same as the actual manipulation order. However, the previous manipulation always hides beneath the subsequent manipulation. Think that it is challenging to solely take out the center of an “onion” without peeling off its skin layers. Similarly, the first manipulation is the most challenging to detect, and this challenge hinders the model from gradually capturing sequential traces.
In this paper, we delve into the above perspectives and propose a Texture-aware and Shape-guided Transformer (TS-Former) to expose sequential DeepFakes (see Fig. 1(d)). Our method introduces four major improvements, each derived from a specific perspective. 1) In the encoder, we develop two branches for capturing different types of manipulation traces. One is the vanilla branch, which is composed of conventional Transformer blocks that capture global relationships. Inspired by the effect of textures in forgery detection Liu et al. (2020); Yang et al. (2021), we design a texture-aware branch to capture the subtle manipulation traces. In this branch, we describe a Diversiform Pixel Difference Attention (DPDA) module, which adaptively integrates various texture extraction operations into the self-attention mechanism, with guidance from the vanilla branch. 2) Then we propose a new Bidirectional Interaction Cross-attention (BICA) module to effectively model correlations among the spatial features from the encoder and sequential manipulation annotations. Unlike standard cross-attention, our module simultaneously processes multiple features and employs a multi-level fusion strategy for more effective correlations. 3) To enhance cross-attention, we propose a Shape-guided Gaussian Mapping strategy to provide the priors of manipulation shapes. Instead of estimating the scale and position, we predict the Gaussian distribution of their manipulated shapes using a Variational Auto-encoder (VAE). 4) Subsequently, we invert the order of the manipulation sequences in the training phase. This simple “trick” notably improves the detection performance as expected. Our method is evaluated on a public sequential DeepFake dataset and compared with several state-of-the-art methods, demonstrating the effectiveness of our method in detecting sequential DeepFakes.
The contribution of this paper is summarized as
  1. 1.
    We introduce a texture-aware branch to capture subtle manipulation traces. To achieve this, we describe a Diversiform Pixel Difference Attention (DPDA) module, which adaptively integrates various texture extraction operations into the self-attention mechanism.
  2. 2.
    We describe a Bidirectional Interaction Cross-attention (BICA) module to effectively model the correlations among spatial and sequential features in a multi-level fusion strategy. Then we propose a Shape-guided Gaussian Mapping strategy to enhance cross-attention.
  3. 3.
    By intriguingly inverting the prediction order of manipulation sequences, we achieve notable performance gains.
  4. 4.
    Through extensive experiments, we demonstrate the efficacy of our method and thoroughly study the effect of each component of our method for further exploration.
2 Related Works
DeepFake Detection. To combat the threats posed by DeepFake, a considerable amount of methods have been proposed for DeepFake detection Li et al. (2018); Yang et al. (2019); Gu et al. (2022); Xu et al. (2023); Dong et al. (2023); Guo et al. (2023); Wang et al. (2023). These methods can typically be classified into different categories in terms of the way of feature capturing. The methods of Li et al. (2018); Yang et al. (2019); Agarwal et al. (2019); Qi et al. (2020); Yu et al. (2024) expose DeepFake by capturing abnormal biological signals, such as eye blinking, behavior patterns, or heart rhythm. Other methods focus on extracting the spatial artifacts introduced either by face generation or the blending operations Li and Lyu (2019); Li et al. (2020); Zhao et al. (2021c); Shiohara and Yamasaki (2022). Instead of capturing the spatial artifacts, the methods of Li et al. (2021); Liu et al. (2021); Dzanic et al. (2020) demonstrate the effectiveness of frequency information and capture representative traces in the frequency domain. Moreover, many recent methods design dedicated models to capture the traces automatically Zhao et al. (2021a); Guo et al. (2023); Wang et al. (2023) However, these methods can only predict authenticity, which can not handle and analyze the multi-step alterations in practice.
Sequential DeepFake Detection. Sequential DeepFake detection is proposed to predict the sequential manipulations in order. This task is generally formulated as the image-to-sequence problem and utilizes sequential-related architectures for detection. The method of Shao et al. (2022) solves this task using a plain Transformer architecture. In this paper, we further explore this task and describe a new method dedicated to sequential DeepFake detection. Our method features four major improvements and has demonstrated effectiveness in experiments.
Figure 2: Overview of the spatial encoder (top). Illustration of DPDA module (bottom). MSA denotes Multi-head Self-attention.
3 Method
Our method consists of a Transformer encoder and a Transformer decoder. The encoder takes the face images as input and extracts spatial manipulation features, while the decoder models the sequential relation based on these features and sequence annotations to predict the facial manipulation sequence.
The encoder is composed of a vanilla branch and a new texture-aware branch. In the texture-aware branch, we introduce the Diversiform Pixel Difference Attention (DPDA) module to capture subtle manipulation traces (Sec. 3.1). For the decoder, we propose a Bidirectional Interaction Cross-attention strategy to fuse the extracted spatial features with the sequence annotation embeddings, enabling effective modeling of sequential relations (Sec. 3.2). To provide more guidance, we introduce Shape-guided Gaussian Mapping to enhance the effect of cross-attention in sequential modeling (Sec. 3.2). We then introduce the procedure for inverting the prediction orders (Sec. 3.2) and outline the overall loss functions (Sec. 3.3).
3.1 Spatial Encoder
Tokenization. Denote a face image as $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$. To tokenize this face image, we employ a simple CNN stem following He et al. (2015). Let a CNN stem be $\mathcal{F}$. The input tokens can be obtained as $\mathbf{x}^{0}=\mathcal{F}(\mathbf{I})+\mathbf{p},\mathbf{z}^{0}\in\mathbb{R}^{h\times w\times c}$, where $\mathcal{F}(\mathbf{I})$ is the feature maps from CNN model and $\mathbf{p}$ is a positional embedding. Then we flatten the spatial dimension of these tokens to one dimension as the input of the encoder.
Vanilla Branch. This branch contains several Transformer blocks and each block calculates a multi-head self-attention. Denote the index of a block as $l$ and the input token of this block as $\mathbf{x}^{l}$. Thus the operation of this block can be written as
$\hat{\mathbf{x}}^{l}$ $=\mathbf{x}^{l}+{\rm MSA}({\rm LN}(\mathbf{x}^{l})),$ (1)
$\hat{\mathbf{x}}^{l+1}$ $=\hat{\mathbf{x}}^{l}+{\rm FFN}({\rm LN}(\hat{\mathbf{x}}^{l})),$
where ${\rm MSA}$ denotes the multi-head self-attention operation, ${\rm LN}$ denotes the LayerNorm Ba et al. (2016), FFN is the feed-forward network, which consists of two FC layers with GELU Hendrycks and Gimpel (2016).
Let $\mathbf{x}$ be the input tokens of MSA layer. In each head, three projection matrices are applied to generate query, key, and value features. Then the self-attention is conducted over the input tokens spatially to learn the coarse-grained traces. Within the $i$-th head, the self-attention can be defined as
$Q_{i}=\mathbf{x}W^{Q}_{i},K_{i}=\mathbf{x}W^{K}_{i},V_{i}=\mathbf{x}W^{V}_{i},$ (2)
$h_{i}={\rm Softmax}(Q_{i}K^{\top}_{i}/\sqrt{d_{i}})V_{i},$
${\rm MSA}(\mathbf{x})={\rm Concat}(h_{1},...,h_{n})W^{o},$
where $W^{Q}_{i},W^{K}_{i},W^{V}_{i}$ are the projection matrices, $d_{i}$ is the number of hidden dimensions, and $h_{i}$ is the self-attention of $i$-th head. Then we concatenate all these heads and perform a linear transformation with $W^{o}$ to obtain the final output.
Texture-aware Branch. To extract these subtle manipulation traces, we develop a new Transformer block designed specifically for extracting texture features. Our block draws inspiration from central difference operators Yu et al. (2020b); Juefei-Xu et al. (2017), which are convolutions used for detecting local textures. In our method, we derive and extend their essence from the convolution style to the self-attention mechanism, and propose a Diversiform Pixel Difference Attention (DPDA) module that captures fine-grained spatial traces while retaining the global view within the vanilla self-attention.
1) Revisit of Central Difference Convolution. Central difference convolution is a special case of convolution, which emphasizes the center-oriented gradients in the local receptive field region. Given a feature map $\mathbf{x}$, this operation can be formulated as
$\mathbf{x}^{\prime}(p_{0})=$ $\sum_{p_{n}\in\mathcal{O}}\mathbf{w}(p_{n})\cdot(\mathbf{x}(p_{0}+p_{n})-\mathbf{x}(p_{0})),$ (3)
where $\mathbf{w}$ is the convolution kernel, $\mathcal{O}$ denotes the local receptive field, $p_{0}$ indicates the current location of input feature map $\mathbf{x}$ and output feature map $\mathbf{x}^{\prime}$, and $p_{n}$ is the relative position in $\mathcal{O}$.
2) Diversiform Pixel Difference Attention. Inspired by the central difference operations, we extend these operations and integrate such spirit into the self-attention mechanism following Yu et al. (2020b, a); Miao et al. (2023). Specifically, given the feature map $\mathbf{x}$, we first perform a vanilla convolution on it to generate the query feature, which is defined as
$\mathbf{x}^{Q}(p_{0})=$ $\sum_{p_{n}\in\mathcal{O}}\mathbf{w}(p_{n})\cdot\mathbf{x}(p_{0}+p_{n}).$ (4)
Then we would like to capture the fine-grained spatial features as the source of key and value features. Since the central difference operations focus on the center-oriented gradients, they may overlook other texture patterns, which hinders the effectiveness of capturing manipulation traces. To obtain a diverse range of gradient information, we leverage the concept of Extended Local Binary Pattern (ELBP), which can encode the pixel relations from various perspectives, including central, angular, and radial directions Su et al. (2021). The angular direction calculates the difference between adjacent pixels in the outline of the kernel in a certain order, while the radial direction calculates the difference between radial adjacent pixels, see Fig. 2. To obtain the key and value features, we adaptively ensemble these operations using a learnable coefficient vector, which is formulated as
$\mathbf{x}^{K,V}(p_{0})=\sum_{p_{n}\in\mathcal{O}}\mathbf{w}(p_{n})\cdot$ $\left(\alpha_{c}\cdot\underbrace{(\mathbf{x}(p_{0}+p_{n})-\mathbf{x}(p_{0}))}_{\mathcal{D}_{c}}\right.$ (5)
$+\alpha_{a}\cdot\underbrace{(\mathbf{x}(p_{0}+p_{n})-\mathbf{x}(p_{n-1}))}_{\mathcal{D}_{a}}$
$+\alpha_{r}\cdot\left.\underbrace{(\mathbf{x}(p_{0}+p_{n})-\mathbf{x}(p_{\phi(n)}))}_{\mathcal{D}_{r}}\right)$
$=\sum_{p_{n}\in\mathcal{O}}\mathbf{w}(p_{n})\cdot$ $\left(\underbrace{\left[\alpha_{c},\alpha_{a},\alpha_{r}\right]}_{\bm{\alpha}}\left[\begin{array}[]{l}\mathcal{D}_{c}\\ \mathcal{D}_{a}\\ \mathcal{D}_{r}\end{array}\right]\right).$
In this equation, $\mathcal{D}_{c},\mathcal{D}_{a},\mathcal{D}_{r}$ denote central, angular, and radial difference operations respectively. We use $3\times 3$ as the local receptive field $\mathcal{O}$ for instance and set $\mathcal{O}=\{(-1,-1),(-1,0),...,(1,1)\}$. Thus $\phi(n)$ in radial difference operation can defined as $\phi(n)=(2x,2y)$ if $n=(x,y)$. $\bm{\alpha}$ is the coefficient vector to balance the contribution of each operation.
Given that different DeepFake face images may contain different manipulation traces, we design a learning scheme to adaptively balance the contribution strength of different DPDA operations. As our first branch concentrates on extracting coarse-grained information, its intermediate features can provide global instruction on the composition of manipulations. Denote the feature map of the corresponding block in our coarse-grained branch as $\mathbf{x}^{\prime}$. To achieve this adaptive learning, we design a sub-network $\mathcal{H}$ that contains several convolution layers. This network learns the $\bm{\alpha}$ based on $\mathbf{x}^{\prime}$ as
$\bm{\alpha}={\rm Softmax}(\mathcal{H}(\mathbf{x}^{\prime})).$ (6)
After obtaining $\mathbf{x}^{Q}$ and $\mathbf{x}^{K,V}$, we project the query feature $\mathbf{x}^{Q}$ into query tokens, and project $\mathbf{x}^{K,V}$ into key and value tokens. Within each head, ${Q}_{i}=\mathbf{x}^{Q}{W}^{Q}_{i},{K}_{i}=\mathbf{x}^{K,V}{W}^{K}_{i},{V}_{i}=\mathbf{x}^{K,V}{W}^{V}_{i}$. Then the multi-head self-attention is calculated using the Eq. (2).
3) Diversiform Pixel Difference Transformer Block. This block integrates the Diversiform Pixel Difference Attention (DPDA) module, which is similar to the Eq. (1). The overall process can be written as
$\hat{\mathbf{x}}^{l}$ $=\mathbf{x}^{l}+{\rm DPDA}({\rm LN}(\mathbf{x}^{l})),$ (7)
$\hat{\mathbf{x}}^{l+1}$ $=\hat{\mathbf{x}}^{l}+{\rm FFN}({\rm LN}(\hat{\mathbf{x}}^{l})),$
With this strategy, this branch can effectively extract subtle manipulation traces, enhancing the performance in sequential detection.
3.2 Sequential Decoder
In order to predict sequential manipulations, we develop a sequential decoder architecture, which takes as input the features from the encoder and the embeddings of manipulation sequences. To effectively integrate these types of features into the decoder, we propose a Bidirectional Interactive Cross-attention (BICA) strategy that can properly integrate the features from the encoder into the decoder. Moreover, to improve the capacity of cross-attention, we introduce a shape-guided Gaussian mapping strategy to enhance its effectiveness.
Spatial and Sequential Feature Fusing. To fully leverage the benefits of both coarse-grained and fine-grained features, we propose a Bidirectional Interaction Cross-attention (BICA) module. This module employs a hierarchical attention mechanism to explore the correlation between features and sequential annotations.
1) Revisit of Conventional Cross-attention The cross-attention is typically employed in the decoder of the transformer for sequence tasks. Unlike self-attention mechanisms, the query, and key and value features usually originate from different sources. Given two input tokens from different sources $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$, the multi-head cross-attention (MCA) can be defined as
$Q_{i}=\mathbf{x}_{1}W^{Q}_{i},K_{i}=\mathbf{x}_{2}W^{K}_{i},V_{i}=\mathbf{x}_{2}W^{V}_{i},$ (8)
$h_{i}={\rm Softmax}(Q_{i}K^{\top}_{i}/\sqrt{d_{i}})V_{i},$
${\rm MCA}(\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{2})={\rm Concat}(h_{1},...,h_{n})W^{o},$
2) Bidirectional Interactive Cross-attention. Let $\mathbf{x}^{c}$ and $\mathbf{x}^{f}$ represent the features from the coarse-grained and fine-grained branches. Let $\mathbf{x}^{s}$ be the embedding of sequential manipulation annotations. In our task, it is intuitive to perform separate MCA operations on $\mathbf{x}^{s}$ with $\mathbf{x}^{c}$ and $\mathbf{x}^{f}$. However, this integration fails to explore deeper connections between these three features. Inspired by the concept of multi-modal cross-attentions Yu et al. (2019); Wei et al. (2020); Zhang et al. (2023), we propose Bidirectional Interactive Cross-attention (BICA) which employs a multi-level fusion strategy to fully excavate the underlying correlations. The overview of this module is shown in Fig. 3.
Figure 3: Overview of the sequential decoder (top). Illustration of Bidirectional Interactive Cross-attention (BICA). MCA denotes Multi-head Cross-attention.
Firstly, we perform MSA on $\mathbf{x}^{s}$ to find the correlations of sequential manipulation annotations, which is then subjected to MCA with $\mathbf{x}^{c}$ and $\mathbf{x}^{f}$ to build the cross-attention. Then these processed features are summed up and applied LayerNorm to generate multi-modal features $\mathbf{a}^{c}$ and $\mathbf{a}^{f}$. These operations can be written as
$\tilde{\mathbf{x}}^{s}$ $={\rm MSA}(\mathbf{x}^{s}),$ (9)
$\tilde{\mathbf{a}}^{c}$ $={\rm MCA}(\tilde{\mathbf{x}}^{s},\mathbf{x}^{c},\mathbf{x}^{c}),\mathbf{a}^{c}={\rm LN}(\tilde{\mathbf{a}}^{c}+\tilde{\mathbf{x}}^{s}),$
$\tilde{\mathbf{a}}^{f}$ $={\rm MCA}(\tilde{\mathbf{x}}^{s},\mathbf{x}^{f},\mathbf{x}^{f}),\mathbf{a}^{f}={\rm LN}(\tilde{\mathbf{a}}^{f}+\tilde{\mathbf{x}}^{s}).$
We then perform interactive cross-attention between features of $\mathbf{a}^{c}$ and $\mathbf{a}^{f}$ in two directions as
$\tilde{\mathbf{a}}^{c,f}$ $={\rm MCA}(\mathbf{a}^{c},\mathbf{a}^{f},\mathbf{a}^{f}),\mathbf{a}^{c,f}={\rm LN}(\tilde{\mathbf{a}}^{c,f}+\mathbf{a}^{c}),$ (10)
$\tilde{\mathbf{a}}^{f,c}$ $={\rm MCA}(\mathbf{a}^{f},\mathbf{a}^{c},\mathbf{a}^{c}),\mathbf{a}^{f,c}={\rm LN}(\tilde{\mathbf{a}}^{f,c}+\mathbf{a}^{f}).$
Next, we fuse the sequential features $\tilde{\mathbf{x}}^{s}$ with the enhanced spatial features $\mathbf{a}^{c,f}$ and $\mathbf{a}^{c,f}$, following the operations as
$\mathbf{A}$ $=[\mathbf{a}^{c},\mathbf{a}^{f},\mathbf{a}^{c,f},\mathbf{a}^{f,c}],$ (11)
$\tilde{\mathbf{x}}^{o}$ $={\rm MCA}(\tilde{\mathbf{x}}^{s},\mathbf{A},\mathbf{A}),$
${\mathbf{x}}^{o}$ $={\rm LN}(\tilde{\mathbf{x}}^{o}+\tilde{\mathbf{x}}^{s}).$
Through this fusion strategy, the sequential and spatial features can be fully interacted with and expose the underlying correlations.
Shape-guided Gaussian Mapping In comparison to general image-sequence tasks, sequential manipulation annotations in our case are relatively short and may not provide sufficient guidance. To address this limitation, we propose a Shape-guided Gaussian Mapping strategy to leverage the strong spatial position priors associated with each manipulation component in the faces. Specifically, we assume that the position of each facial manipulation follows a Gaussian distribution. By learning these distributions, we can determine the probability of each position on a face image being manipulated. To simulate these Gaussian distributions, we construct a Variational Auto-encoder (VAE) Kingma and Welling (2013) with the image $\mathbf{I}$ as input. The VAE architecture contains an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$. The encoder transforms the input feature into means $\bm{\mu}$ and standards $\bm{\sigma}$ of these Gaussian distributions, while the decoder reconstructs the input images from the sampled latent code. Through the reconstruction process, the encoder learns to capture the spatial manipulation information and formulate them as Gaussian distributions. Next, we create a grid map $M$ that is composed of the $x$- and $y$- coordinates of each position. This grid map is then used to predict the probability of each position in different facial manipulations. These operations can be formulated as
$\bm{\mu},\bm{\sigma}=\mathcal{E}(\mathbf{I}),$ (12)
$S=\mathcal{N}(M\;|\;\bm{\mu},\bm{\sigma}),$
where $S$ is the probability map for each spatial position. This map is integrated into the MCA operations ${\rm MCA}(\tilde{\mathbf{x}}^{s},\mathbf{x}^{c},\mathbf{x}^{c})$ and ${\rm MCA}(\tilde{\mathbf{x}}^{s},\mathbf{x}^{f},\mathbf{x}^{f})$ in Eq. (9) respectively. We take the integration in ${\rm MCA}(\tilde{\mathbf{x}}^{s},\mathbf{x}^{c},\mathbf{x}^{c})$ for example as
$Q_{i}=\tilde{\mathbf{x}}^{s}W^{Q}_{i},K_{i}=\mathbf{x}^{c}W^{K}_{i},V_{i}=\mathbf{x}^{c}W^{V}_{i},$ (13)
$h_{i}={\rm Softmax}(Q_{i}K^{\top}_{i}/\sqrt{d_{i}}+\eta\cdot S)V_{i},$
${\rm MCA}(\tilde{\mathbf{x}}^{s},\mathbf{x}^{c},\mathbf{x}^{c})={\rm Concat}(h_{1},...,h_{n})W^{o}.$
By inserting the shape priors of facial manipulation, the cross-attention is greatly enhanced to capture the spatial and sequential forgery traces.
Figure 4: The impact of the sequential manipulation on facial attributes. The second manipulation impacts the first manipulation on (left) eye and (right) eyeglasses.
Inverted Order Prediction. Based on our analysis, we invert the annotation of the manipulation sequence for training. For instance, given a manipulated face with annotation of “eyes-nose-eyebrow-lips-hair”, we reorder it as “hair-lips-eyebrow-nose-eyes”, and convert it to $\mathbf{x}^{s}$ for sequential modeling. By doing so, the model first predicts the last manipulation of “hair”. By isolating the disturbance caused by “hair”, the model can then focus on predicting the next manipulation, which is “lips”. The rationale behind this strategy is that the manipulation of the nose can affect the manipulation of the eyes, making the detection of eye manipulations more challenging. On the other hand, the manipulation in hair is not influenced by other manipulations, making it more easily detectable (see Fig. 4). The advantage of this strategy is that, compared to forward sequence predictions, each step in the inverted sequence predictions is only influenced by itself and is independent of other types of manipulations.
3.3 Loss Function
In the training phase, we employ an auto-regressive mechanism Graves (2013) to model sequence relationships with the cross-entropy loss $L_{\rm ce}$. This loss measures the error of predicted manipulation sequence and ground truth. For the shape-guided module, we perform the general image reconstruction loss $L_{\rm rec}$ and KLD loss $L_{\rm kld}$ used in VAE Kingma and Welling (2013). The overall loss function is expressed as
$L=L_{\rm ce}+\lambda_{1}L_{\rm rec}+\lambda_{2}L_{\rm kld},$ (14)
where ${\lambda}_{1},{\lambda}_{2}$ are weights for balancing these loss terms.
Table 1: Performance of different methods on detecting sequential DeepFakes.
Methods Facial-components Facial-attributes Average
Fixed-Acc Adaptive-Acc Fixed-Acc Adaptive-Acc Fixed-Acc Adaptive-Acc
Multi-Cls (ResNet-34) 69.66 50.54 66.99 46.68 68.33 48.61
Multi-Cls (ResNet-50) 69.65 50.57 66.66 46.00 68.16 48.29
DETR (ResNet-34) 69.87 50.63 67.93 48.15 68.90 49.39
DETR (ResNet-50) 69.75 49.84 67.62 47.99 68.69 48.92
DRN (ResNet-34) 66.06 45.79 64.42 43.20 65.24 44.50
MA (ResNet-34) 71.31 52.94 67.58 47.48 66.45 50.21
Two-Stream (ResNet-34) 71.92 53.89 66.77 46.38 69.35 50.14
SeqFake-Former (ResNet-34) 70.88 52.96 67.85 48.36 69.37 50.66
SeqFake-Former (ResNet-50) 69.41 50.22 68.38 48.97 68.90 49.60
TS-Former (Ours) (ResNet-34) 74.70 58.43 69.13 49.68 71.92 54.06
TS-Former (Ours) (ResNet-50) 73.32 56.08 68.86 49.67 71.09 52.88
Table 2: Ablation study on the facial-components datasets.
Setting Resnet-34 Resnet-50
Fixed-Acc Adaptive-Acc Fixed-Acc Adaptive-ACC
Baseline 69.65 (+0.00) 50.43 (+0.00) 68.84 (+0.00) 49.40 (+0.00)
Baseline+DPDA 70.24 (+0.59) 51.87 (+1.44) 69.23 (+0.39) 51.05 (+1.65)
Baseline+DPDA+BICA 71.00 (+1.35) 52.67 (+2.24) 71.55 (+2.71) 53.06 (+3.66)
Baseline+DPDA+BICA+IOP 73.32 (+3.67) 56.65 (+6.22) 72.65 (+3.81) 55.03 (+5.63)
Baseline+DPDA+BICA+IOP+SGM 74.70 (+5.05) 58.43 (+8.00) 73.32 (+4.48) 56.08 (+6.68)
4 Experiments
4.1 Experimental Settings
Datasets. Our experiments are conducted on Sequential DeepFake Dataset (SDD) Shao et al. (2022). This dataset contains two types of sequential DeepFake manipulations, the facial components manipulation and the facial attributes manipulation. The first track comprises $35,166$ face images that are subjected to replacing operations on specific facial components. There are $28$ manipulation types in this track, with a maximum sequence length of five. The second track contains $49,920$ face images with $26$ types of attribute manipulations. The maximum length of each manipulation is also five. Its split of training, validating, and testing set is $8:1:1$.
Evaluation Metrics. In line with previous work Shao et al. (2022), we employ Fixed Accuracy (Fixed-Acc) and Adaptive Accuracy (Adaptive-Acc) for evaluation. To calculate the Fixed-Acc score, the sequence annotations shorter than five are padded with the label “no manipulation” and then compare the predictions with the padded sequence annotation. On the other hand, due to the nature of the auto-regress mechanism, the model can stop predicting upon encountering the End-of-Sequence (EOS) symbol, resulting in various lengths of the predicted sequence. The adaptive-Acc score is calculated by comparing the predictions with the sequence annotations in terms of dynamic lengths.
Implementation Details. We utilize ResNet-34 He et al. (2015) as our CNN stem in main experiment. Each branch in the spatial encoder has two Transformer blocks. Each block contains four attention heads. The sequential decoder consists of two Transformer blocks. In the training phase, the image size is set to $256\times 256$ with augmentations such as horizontal Flipping and adjusting the brightness of the image. We employ the AdamW optimizer with a weight decay of $10^{-4}$. The initial learning rates are set to $10^{-4}$ for CNN stem and $10^{-4}$ for Transformer architecture. We utilize a warm-up strategy of $20$ epochs, with a batch size of $32$. The total epochs are set to $170$, with a $10\%$ reduction in learning rate every $50$ epochs. The parameters in the loss function are set as: $\lambda_{1}=0.1,\lambda_{2}=0.01$.
4.2 Results
Table 1 shows the performance of our method compared to other counterparts, including Multi-Cls, DETR Carion et al. (2020), DRN Wang et al. (2019), MA Zhao et al. (2021b), Two-Stream Luo et al. (2021), SeqFake-Former Shao et al. (2022) respectively. Multi-Cls and DETR serve as baseline methods adapted from models for general vision tasks. Multi-Cls performs multi-label classification by directly classifying the manipulated images into multiple classes. DETR is a modified object detector that replaces object queries with manipulation annotation. DRN, MA, and Two-Stream are DeepFake detection methods adapted by substituting the binary classification head with a multi-classification head. These methods are all referrenced from Shao et al. (2022). SeqFake-Former is a recent dedicated method for sequential DeepFake detection. We reproduce this method rigorously using their release codes. As shown in Table 1, our method outperforms baseline methods by a large margin, averaging $71.92\%$ and $54.06\%$ in Fixed-Acc and Adaptive-Acc scores, respectively. Moreover, our method surpasses the DeepFake detection methods in both tracks. Notably, compared to SeqFakeFormer, our method achieves improvements by $3.82\%,5.47\%$ and $1.28\%,1.32\%$ in these two tracks.
The results also indicate that all methods perform better in Fixed-Acc than in Adaptive-Acc. This observation is consistent with the findings in Shao et al. (2022) that predicting manipulation sequences with varying lengths is a more challenging task. In real-world applications, manipulation sequences with varying lengths are more practical, thus the improvement of our method in this metric sufficiently represents the practicability of the proposed modules.
Intriguingly, we observe that ResNet-50 performs slightly worse than ResNet-34 across all methods. This contrasts with the expectation that a larger model should have stronger capacity. We discuss this phenomenon in Sec. 4.4.
4.3 Ablation Study
Our method consists of several key components, including the Diversified Pixel Difference Attention module (DPDA), Bidirectional Interactive Cross-attention (BICA), Inverted Order Prediction (IOP), and Shape-guided Gaussian Mapping (SGM). To evaluate the effectiveness of each component, we conduct a series of ablation studies on the facial-components track using the stem of ResNet-34 and ResNet-50.
Specifically, we designed different models based on the following settings: 1) A baseline model with a Transformer architecture that only consists of basic encoders and decoders. 2) Adding the DPDA module to the baseline model, which generates two spatial features from the encoder. Then we directly add these two features and feed them into the sequential decoder. 3) Adopting the BICA strategy on the baseline model for feature integration. 4) Incorporating inverted order prediction. 5) Inserting the SGM module into BICA strategy.
The results of the ablation studies are shown in Table 2. It can be observed that each component contributes to the overall performance. Taking Resnet-34 as an example. The DPDA module improves the Fixed-Acc and Adaptive-Acc by $0.59\%$ and $1.44\%$, respectively. The BICA strategy further enhances the performance by $0.76\%$ and $0.8\%$ in Fixed-Acc and Adaptive-Acc. After applying IOP, the performance is significantly improved by $2.32\%$ in Fixed-Acc and $3.98\%$ in Adaptive-Acc, showing the effectiveness of this strategy in gradually capturing manipulation traces. Finally, the integration of SGM yields further performance, achieving the best performance when all components are applied.
Figure 5: The visualization of our model and SeqFakeFormer under the (a) eyebrow-eye-hair-nose-lip and (b) nose-eyebrow-lip-eye-hair manipulating annotations.
Table 3: Performance on Different CNN Stems.
Stem Fixed-Acc Adaptive-Acc
SFF Ours SFF Ours
ResNet-18 68.41 69.97 49.55 52.26
ResNet-34 70.88 74.70 52.96 58.43
Resnet-50 69.41 73.32 50.22 56.08
Resnet-101 68.87 73.46 49.06 56.77
4.4 Discussion
Attention Visualization. To assess the learning ability of our method, we visualize the attention maps of our method compared with Seq-FakeFormer under two different manipulation annotations, eyebrow-eye-hair-nose-lip and nose-eyebrow-lip-eye-hair. Fig. 5 shows the Grad-CAM Selvaraju et al. (2019) of these methods. We observe that under the same annotation, our method can identify the manipulated regions more precisely. In contrast, Seq-FakeFormer can localize the major manipulated regions but may overlook small manipulations with subtle traces. This result highlights the effectiveness of our method in capturing sequential manipulation traces.
Furthermore, we also observe that our method can highlight regions that are not annotated as manipulation. We attribute this to the imperfect nature of GAN in the feature space, causing subtle influences on some neighboring areas during the multi-step facial manipulation process. Our method can capture this subtle alternation as evidence for exposing the sequential manipulations.
Controversy Regarding CNN Stem. From Table 3 we observe an unexpected phenomenon where using ResNet-34 as CNN stem achieves better performance than ResNet-50. To further investigate this, we validate our method using ResNet-18 and ResNet-101. Table 3 shows the performance comparison of our method and SeqFake-Former (SFF) on these architectures. It can be seen that the performance of ResNet-18 is notably degraded compared to ResNet-34, aligning with our expectations. However, when using ResNet-101 as the Stem, it yields lower performance than ResNet-50. This finding contradicts the general understanding that increasing the model size and parameters can improve feature extraction capacity. One possible explanation for this phenomenon is that the length of the manipulation sequence is relatively short (less than five), which increases the likelihood of overfitting in large models. This suggests the need to strike a balance between model size and performance in the future.
5 Conclusion
We introduce a novel Texture-aware and Shape-guided Transformer to enhance detection performance. Compared to the classic Transformer-based methods, we introduce four improvements. Firstly, we describe a new Diversiform Pixel Difference Attention module (DPDA) to capture subtle manipulation traces. Moreover, we introduce a Bidirectional Interaction Cross-attention module (BICA) to explore deep correlations between spatial and sequential features. To further enhance the cross-attention, we describe a Shape-guided Gaussian mapping strategy to provide the priors of manipulation shapes. Lastly, we improve the performance by inverting the prediction order from forward to backward. The results obtained on public datasets demonstrate the efficacy of our method in sequential DeepFake detection.

### v4

@@
-Texture-aware and Shape-guided Transformer for Sequential DeepFake Detection
+Texture, Shape and Order Matter: A New Transformer Design for Sequential DeepFake Detection
 Abstract
-Sequential DeepFake detection is an emerging task that aims to predict the manipulation sequence in order.
-Existing methods typically formulate it as an image-to-sequence problem, employing conventional Transformer architectures for detection.
+Sequential DeepFake detection is an emerging task that predicts the manipulation sequence in order.
+Existing methods typically formulate it as an image-to-sequence problem, employing conventional Transformer architectures.
 However, these methods lack dedicated design and consequently result in limited performance.
-In this paper, we propose a novel Texture-aware and Shape-guided Transformer to enhance detection performance.
-Our method features four major improvements.
-Firstly, we describe a texture-aware branch that effectively captures subtle manipulation traces with the Diversiform Pixel Difference Attention module.
-Then we introduce a Bidirectional Interaction Cross-attention module that seeks deep correlations among spatial and sequential features, enabling effective modeling of complex manipulation traces.
-To further enhance the cross-attention, we describe a Shape-guided Gaussian mapping strategy, providing initial priors of the manipulation shape.
-Finally, observing that the latter manipulation in a sequence may influence traces left in the earlier one, we intriguingly invert the prediction order from forward to backward, leading to notable gains as expected.
+As such, this paper describes a new Transformer design, called TSOM, by exploring three perspectives: Texture, Shape, and Order of Manipulations.
+Our method features four major improvements: ❶ we describe a new texture-aware branch that effectively captures subtle manipulation traces with a Diversiform Pixel Difference Attention module.
+❷ Then we introduce a Multi-source Cross-attention module to seek deep correlations among spatial and sequential features, enabling effective modeling of complex manipulation traces.
+❸ To further enhance the cross-attention, we describe a Shape-guided Gaussian mapping strategy, providing initial priors of the manipulation shape.
+❹ Finally, observing that the subsequent manipulation in a sequence may influence traces left in the preceding one, we intriguingly invert the prediction order from forward to backward, leading to notable gains as expected.
 Extensive experimental results demonstrate that our method outperforms others by a large margin, highlighting the superiority of our method.
@@
 One such technique, DeepFake, allows for effortless manipulation of faces, with results that are often imperceptible to the naked eye.
-This technique has found applications in virtual reality and movie special effects, but it also raises severe security concerns He et al.
-(2021).
-To address these concerns, various forensic methods have been proposed to expose DeepFakes, garnering considerable attention.
-Conventional forensics methods mainly focus on determining whether a given face image is authentic or forged, which typically falls into the binary classification task.
+This technique has found applications in virtual reality and movie special effects, but it also raises severe security concerns [9].
+To address these concerns, various forensic methods have been proposed to expose DeepFakes, garnering considerable attention [15, 34, 14].
+Conventional forensics methods mainly focus on determining whether a given face image is authentic or forged, which typically falls into the binary classification task (see Fig.
+1(a)).
 However, with the rapid evolution of face editing tools, the manipulation process has become more versatile and user-friendly, enabling the use of multiple specific operations iteratively until the desired effects are reached.
-Under this circumstance, existing methods have limitations as they can only identify the final step of manipulation, providing limited insights into the manipulation details and hindering traceability and interpretability.
-To overcome this issue, a new forensics task called Sequential DeepFake Detection Shao et al.
-(2022) has emerged recently, aiming to uncover the sequences of face manipulations (see Fig.
-1(a,b)).
-This task expands the scope of DeepFake detection and poses a new challenge to this field.
-Figure 1: (a) and (b) describe the difference between conventional DeepFake detection and sequential DeepFake detection.
-(c) illustrates the process of classic sequential DeepFake detection methods, while (d) represents our proposed method.
-Typically, Transformer-based architectures are employed in sequential DeepFake detection to capture traces of different manipulations and predict the manipulation sequences, following the pipeline of image-to-sequence Vinyals et al.
-(2015).
-These methods commonly employ a plain Transformer encoder to extract spatial clues and utilize a Transformer decoder to capture the sequential clues (see Fig.
+In this context, existing methods can only identify the final step of manipulation, providing limited insights into the manipulation details.
+To overcome this limitation, a new forensics task called Sequential DeepFake Detection (SDD) [23] has emerged recently.
+In contrast to the conventional DeepFake detection (i.e., One-step DeepFake Detection (ODD)), SDD attempts to uncover the sequences of face manipulations in order.
+For example, users might create a DeepFake face by sequentially editing regions like hair, eyes, and lips.
+Given such a DeepFake face, SDD can predict the manipulation sequence as “Hair-Eye-Lip” (see Fig.
+1(b)).
+This emerging task improves the practicality of DeepFake detection in real-world scenarios, offering more comprehensive evidence for traceability and interoperability (e.g., pinpointing where, when, and what manipulations are applied).
+This endeavor expands the horizon of DeepFake detection and poses a fresh challenge to the field.
+Figure 1: (a) and (b) correspond to One-step DeepFake Detection (ODD) and sequential DeepFake Detection (SDD).
+(c) illustrates the process of typical Sequential DeepFake Detection (SDD) architecture.
+(d) represents our proposed method TSOM.
+In contrast to (c), our method focuses on exploring three perspectives: Texture, Shape, and Order of Manipulations, and introduces four major improvements: a Diversiform Pixel Difference Attention (DPDA) module, a Multi-source Cross-attention (MSCA) module, a Shape-guided Gaussian Mapping (SGM) strategy and an Inverted Order Prediction (IOP).
+In general, the task of SDD can be conceptualized as an image-to-sequence problem [27], where the typical solution involves leveraging Transformer-based architectures to predict the manipulation sequences [31, 23].
+Concretely, a spatial encoder with conventional self-attention mechanisms is utilized to extract spatial manipulation traces.
+Then, a sequential decoder is employed to establish conventional cross-attention [26] with the encoder features and manipulation annotation sequence, facilitating the capture of sequential manipulation traces (see Fig.
 1(c)).
-Despite these methods showing a certain effectiveness, many critical perspectives are lacking in further exploration.
-  1. $\clubsuit$
-    Fine-grained Traces Extraction: Generative models often exhibit imperfect semantic disentanglement.
-One specific manipulation may affect the entire facial structure and leave subtle traces on unintended manipulated regions.
-To expose the sequential manipulation, existing methods usually utilize a general Transformer encoder to capture those traces.
-However, this encoder concentrates on the global relationship of the entire faces and may fail to capture the subtle manipulation traces.
-  2. $\clubsuit$
-    Manipulation Shape Guidance: In the training stage, existing methods enhance the modeling of sequential relation in the decoder by estimating the scale and position of manipulation regions as priors.
-However, such priors are impractical as the manipulation regions are likely irregular, which could not be well described by only using the scale and position.
-  3. $\clubsuit$
-    Prediction Order: Existing methods predict the manipulation sequences in a forward order, which is the same as the actual manipulation order.
-However, the previous manipulation always hides beneath the subsequent manipulation.
-Think that it is challenging to solely take out the center of an “onion” without peeling off its skin layers.
-Similarly, the first manipulation is the most challenging to detect, and this challenge hinders the model from gradually capturing sequential traces.
-In this paper, we delve into the above perspectives and propose a Texture-aware and Shape-guided Transformer (TS-Former) to expose sequential DeepFakes (see Fig.
+Despite these methods showing certain effectiveness, many critical perspectives are lacking in further exploration: 1) Neglecting texture details: Generative models often exhibit imperfect semantic disentanglement, wherein one specific manipulation may affect the entire facial structure, leaving subtle traces on unintended manipulated regions.
+These subtle traces often manifest more prominently in the textures [17, 33].
+However, the existing methods rely on conventional self-attention mechanisms that concentrate more on global correlations, thus overlooking fine-grained features like textures.
+2) Lack of manipulation shape guidance: The existing pipeline accomplishes detection from a holistic perspective, directly predicting the manipulations without considering their specific shapes.
+Note that the manipulation regions often exhibit various shapes and irregularities.
+Incorporating knowledge about manipulation shapes can guide the network to gain a deeper understanding of the manipulations.
+While existing method [23] incorporates the scale and position of manipulation regions, they could not effectively characterize irregular manipulation regions.
+3) Improper prediction order: Existing methods predict the manipulation sequences in a forward order, the same as the actual manipulation order.
+However, the subsequent manipulation usually affects the appearance of the preceding ones (see Fig.
+2)), i.e., the preceding manipulation “hides beneath” the subsequent one, akin to the layers of an “onion”.
+The first manipulation step (e.g., Hair as shown in Fig.
+1(b)) acts as the core of the onion.
+Extracting this core without peeling away its outer layers (e.g., Eye and lip as shown in Fig.
+1(b)) would be improper.
+Figure 2: The impact of the sequential manipulation on facial attributes. The second manipulation impacts the first manipulation: (left) Eye and (right) Eyeglasses.
+Delving into these perspectives, we propose a new Transformer called TSOM (Texture, Shape and Order of Manipulations) to expose sequential DeepFakes (see Fig.
 1(d)).
-Our method introduces four major improvements, each derived from a specific perspective.
-1) In the encoder, we develop two branches for capturing different types of manipulation traces.
-One is the vanilla branch, which is composed of conventional Transformer blocks that capture global relationships.
-Inspired by the effect of textures in forgery detection Liu et al.
-(2020); Yang et al.
-(2021), we design a texture-aware branch to capture the subtle manipulation traces.
-In this branch, we describe a Diversiform Pixel Difference Attention (DPDA) module, which adaptively integrates various texture extraction operations into the self-attention mechanism, with guidance from the vanilla branch.
-2) Then we propose a new Bidirectional Interaction Cross-attention (BICA) module to effectively model correlations among the spatial features from the encoder and sequential manipulation annotations.
-Unlike standard cross-attention, our module simultaneously processes multiple features and employs a multi-level fusion strategy for more effective correlations.
-3) To enhance cross-attention, we propose a Shape-guided Gaussian Mapping strategy to provide the priors of manipulation shapes.
-Instead of estimating the scale and position, we predict the Gaussian distribution of their manipulated shapes using a Variational Auto-encoder (VAE).
-4) Subsequently, we invert the order of the manipulation sequences in the training phase.
-This simple “trick” notably improves the detection performance as expected.
-Our method is evaluated on a public sequential DeepFake dataset and compared with several state-of-the-art methods, demonstrating the effectiveness of our method in detecting sequential DeepFakes.
-The contribution of this paper is summarized as
-  1. 1.
-    We introduce a texture-aware branch to capture subtle manipulation traces.
-To achieve this, we describe a Diversiform Pixel Difference Attention (DPDA) module, which adaptively integrates various texture extraction operations into the self-attention mechanism.
-  2. 2.
-    We describe a Bidirectional Interaction Cross-attention (BICA) module to effectively model the correlations among spatial and sequential features in a multi-level fusion strategy.
-Then we propose a Shape-guided Gaussian Mapping strategy to enhance cross-attention.
-  3. 3.
-    By intriguingly inverting the prediction order of manipulation sequences, we achieve notable performance gains.
-  4. 4.
-    Through extensive experiments, we demonstrate the efficacy of our method and thoroughly study the effect of each component of our method for further exploration.
+Our method introduces four major improvements: ❶ In the encoder, we develop two branches for capturing different types of manipulation traces.
+One branch, termed the vanilla branch, employs conventional Transformer blocks to capture global relationships.
+The other branch, a new texture-aware branch, is designed to capture subtle manipulation traces.
+Within this branch, we describe a Diversiform Pixel Difference Attention (DPDA) module, which dynamically integrates various texture extraction operations into the self-attention mechanism, instructed by the knowledge from the vanilla branch.
+❷ Then we propose a new Multi-source Cross-attention (MSCA) module to effectively model correlations between spatial features from the encoder and sequential manipulation annotations.
+Unlike traditional cross-attention, our module processes multiple features simultaneously and employs a multi-level fusion strategy for more effective correlations.
+❸ To further enhance cross-attention, we propose a Shape-guided Gaussian Mapping (SGM) strategy to provide priors on manipulation shapes.
+These priors are estimated using a Variational Auto-encoder (VAE) [11].
+❹ During training, we invert the order of the manipulation sequences.
+This simple “trick” notably improves the detection performance, as expected.
+Our method is evaluated on a sequential DeepFake dataset [23] and compared with several state-of-the-art methods.
+The results demonstrate its superiority in detecting sequential DeepFakes.
+Furthermore, we conduct a comprehensive analysis of each component within our method, corroborating the efficacy of the proposed strategies.
 2 Related Works
-DeepFake Detection.
-To combat the threats posed by DeepFake, a considerable amount of methods have been proposed for DeepFake detection Li et al.
-(2018); Yang et al.
-(2019); Gu et al.
-(2022); Xu et al.
-(2023); Dong et al.
-(2023); Guo et al.
-(2023); Wang et al.
-(2023).
+One-step DeepFake Detection (ODD).
+To combat the threats posed by DeepFake, a considerable amount of methods have been proposed for DeepFake detection [14, 34, 6, 32, 3, 7, 30].
 These methods can typically be classified into different categories in terms of the way of feature capturing.
-The methods of Li et al.
-(2018); Yang et al.
-(2019); Agarwal et al.
-(2019); Qi et al.
-(2020); Yu et al.
-(2024) expose DeepFake by capturing abnormal biological signals, such as eye blinking, behavior patterns, or heart rhythm.
-Other methods focus on extracting the spatial artifacts introduced either by face generation or the blending operations Li and Lyu (2019); Li et al.
-(2020); Zhao et al.
-(2021c); Shiohara and Yamasaki (2022).
-Instead of capturing the spatial artifacts, the methods of Li et al.
-(2021); Liu et al.
-(2021); Dzanic et al.
-(2020) demonstrate the effectiveness of frequency information and capture representative traces in the frequency domain.
-Moreover, many recent methods design dedicated models to capture the traces automatically Zhao et al.
-(2021a); Guo et al.
-(2023); Wang et al.
-(2023) However, these methods can only predict authenticity, which can not handle and analyze the multi-step alterations in practice.
-Sequential DeepFake Detection.
-Sequential DeepFake detection is proposed to predict the sequential manipulations in order.
-This task is generally formulated as the image-to-sequence problem and utilizes sequential-related architectures for detection.
-The method of Shao et al.
-(2022) solves this task using a plain Transformer architecture.
-In this paper, we further explore this task and describe a new method dedicated to sequential DeepFake detection.
-Our method features four major improvements and has demonstrated effectiveness in experiments.
-Figure 2: Overview of the spatial encoder (top). Illustration of DPDA module (bottom). MSA denotes Multi-head Self-attention.
+The methods of [14, 34, 1, 20, 35] expose DeepFake by capturing abnormal biological signals, such as eye blinking, behavior patterns, or heart rhythm.
+Other methods focus on extracting the spatial artifacts introduced either by face generation or the blending operations [15, 13, 40, 24].
+Instead of capturing the spatial artifacts, the methods of [12, 16, 4] demonstrate the effectiveness of frequency information and capture representative traces in the frequency domain.
+Moreover, many recent methods design dedicated models to capture the traces automatically [38, 7, 30].
+However, these methods can only predict authenticity, which can not handle and analyze the multi-step alterations in practice.
+Sequential DeepFake Detection (SDD).
+Sequential DeepFake detection is an emerging task focused on predicting the sequential manipulations in order.
+This task is typically formulated as the image-to-sequence problem, utilizing architectures tailored for sequential tasks in detection.
+The work of [23] is pioneering in addressing this problem, employing the Transformer architecture [26], which includes a conventional encoder and decoder with vanilla self-attention mechanisms.
+Subsequently, MMNet [31] improved its performance through the incorporation of multi-collaboration and multi-supervision modules.
+Despite these promising results, this task remains largely unexplored.
 3 Method
-Our method consists of a Transformer encoder and a Transformer decoder.
-The encoder takes the face images as input and extracts spatial manipulation features, while the decoder models the sequential relation based on these features and sequence annotations to predict the facial manipulation sequence.
-The encoder is composed of a vanilla branch and a new texture-aware branch.
+Our method is developed on a Transformer-based architecture comprising a spatial encoder and a sequential decoder.
+The encoder is designed to extract spatial manipulation features, including a vanilla branch and a texture-aware branch.
+The vanilla branch employs several conventional Transformer blocks to capture the global correlations.
 In the texture-aware branch, we introduce the Diversiform Pixel Difference Attention (DPDA) module to capture subtle manipulation traces (Sec.
 3.1).
-For the decoder, we propose a Bidirectional Interaction Cross-attention strategy to fuse the extracted spatial features with the sequence annotation embeddings, enabling effective modeling of sequential relations (Sec.
+The decoder models the sequential relation based on the spatial features and corresponding sequence annotations to predict the facial manipulation sequence.
+In this decoder, we propose a Multi-source Cross-attention (MSCA) strategy to fuse the extracted spatial features with the sequence annotation embeddings, enabling effective modeling of sequential relations (Sec.
 3.2).
-To provide more guidance, we introduce Shape-guided Gaussian Mapping to enhance the effect of cross-attention in sequential modeling (Sec.
-3.2).
-We then introduce the procedure for inverting the prediction orders (Sec.
-3.2) and outline the overall loss functions (Sec.
+Moreover, we introduce Shape-guided Gaussian Mapping (SGM) to enhance the effect of cross-attention in sequential modeling (Sec.
 3.3).
-3.1 Spatial Encoder
-Tokenization.
-Denote a face image as $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$.
-To tokenize this face image, we employ a simple CNN stem following He et al.
-(2015).
-Let a CNN stem be $\mathcal{F}$.
-The input tokens can be obtained as $\mathbf{x}^{0}=\mathcal{F}(\mathbf{I})+\mathbf{p},\mathbf{z}^{0}\in\mathbb{R}^{h\times w\times c}$, where $\mathcal{F}(\mathbf{I})$ is the feature maps from CNN model and $\mathbf{p}$ is a positional embedding.
-Then we flatten the spatial dimension of these tokens to one dimension as the input of the encoder.
-Vanilla Branch.
-This branch contains several Transformer blocks and each block calculates a multi-head self-attention.
-Denote the index of a block as $l$ and the input token of this block as $\mathbf{x}^{l}$.
-Thus the operation of this block can be written as
-$\hat{\mathbf{x}}^{l}$ $=\mathbf{x}^{l}+{\rm MSA}({\rm LN}(\mathbf{x}^{l})),$ (1)
-$\hat{\mathbf{x}}^{l+1}$ $=\hat{\mathbf{x}}^{l}+{\rm FFN}({\rm LN}(\hat{\mathbf{x}}^{l})),$
-where ${\rm MSA}$ denotes the multi-head self-attention operation, ${\rm LN}$ denotes the LayerNorm Ba et al.
-(2016), FFN is the feed-forward network, which consists of two FC layers with GELU Hendrycks and Gimpel (2016).
-Let $\mathbf{x}$ be the input tokens of MSA layer.
-In each head, three projection matrices are applied to generate query, key, and value features.
-Then the self-attention is conducted over the input tokens spatially to learn the coarse-grained traces.
-Within the $i$-th head, the self-attention can be defined as
-$Q_{i}=\mathbf{x}W^{Q}_{i},K_{i}=\mathbf{x}W^{K}_{i},V_{i}=\mathbf{x}W^{V}_{i},$ (2)
-$h_{i}={\rm Softmax}(Q_{i}K^{\top}_{i}/\sqrt{d_{i}})V_{i},$
-${\rm MSA}(\mathbf{x})={\rm Concat}(h_{1},...,h_{n})W^{o},$
-where $W^{Q}_{i},W^{K}_{i},W^{V}_{i}$ are the projection matrices, $d_{i}$ is the number of hidden dimensions, and $h_{i}$ is the self-attention of $i$-th head.
-Then we concatenate all these heads and perform a linear transformation with $W^{o}$ to obtain the final output.
-Texture-aware Branch.
-To extract these subtle manipulation traces, we develop a new Transformer block designed specifically for extracting texture features.
-Our block draws inspiration from central difference operators Yu et al.
-(2020b); Juefei-Xu et al.
-(2017), which are convolutions used for detecting local textures.
+We then introduce the procedure for Inverted Order Prediction (IOP) (Sec.
+3.4) and outline the overall loss functions (Sec.
+3.5).
+3.1 Texture-aware Branch
+[…]
+To extract these subtle manipulation traces, we develop a new Transformer Block for extracting texture features.
+Our block draws inspiration from central difference operators [37, 10], which are convolutions used for detecting local textures.
 In our method, we derive and extend their essence from the convolution style to the self-attention mechanism, and propose a Diversiform Pixel Difference Attention (DPDA) module that captures fine-grained spatial traces while retaining the global view within the vanilla self-attention.
-1) Revisit of Central Difference Convolution.
+Figure 3: Overview of the spatial encoder (top). Illustration of DPDA module (bottom). MSA denotes Multi-head Self-attention.
+Figure 4: Overview of the sequential decoder (left). Illustration of Multi-source Cross-attention(MSCA). MCA denotes Multi-head Cross-attention.
+Revisit of Central Difference Convolution.
 Central difference convolution is a special case of convolution, which emphasizes the center-oriented gradients in the local receptive field region.
 Given a feature map $\mathbf{x}$, this operation can be formulated as
-$\mathbf{x}^{\prime}(p_{0})=$ $\sum_{p_{n}\in\mathcal{O}}\mathbf{w}(p_{n})\cdot(\mathbf{x}(p_{0}+p_{n})-\mathbf{x}(p_{0})),$ (3)
+$\mathbf{x}^{\prime}(p_{0})=$ $\sum_{p_{n}\in\mathcal{O}}\mathbf{w}(p_{n})(\mathbf{x}(p_{0}+p_{n})-\mathbf{x}(p_{0}))$ (1)
 where $\mathbf{w}$ is the convolution kernel, $\mathcal{O}$ denotes the local receptive field, $p_{0}$ indicates the current location of input feature map $\mathbf{x}$ and output feature map $\mathbf{x}^{\prime}$, and $p_{n}$ is the relative position in $\mathcal{O}$.
-2) Diversiform Pixel Difference Attention.
-Inspired by the central difference operations, we extend these operations and integrate such spirit into the self-attention mechanism following Yu et al.
-(2020b, a); Miao et al.
-(2023).
+Diversiform Pixel Difference Attention (DPDA).
+Inspired by the central difference operations, we extend these operations and integrate such spirit into the self-attention mechanism following [37, 36, 19].
 Specifically, given the feature map $\mathbf{x}$, we first perform a vanilla convolution on it to generate the query feature, which is defined as
-$\mathbf{x}^{Q}(p_{0})=$ $\sum_{p_{n}\in\mathcal{O}}\mathbf{w}(p_{n})\cdot\mathbf{x}(p_{0}+p_{n}).$ (4)
+$\mathbf{x}^{Q}(p_{0})=$ $\sum_{p_{n}\in\mathcal{O}}\mathbf{w}(p_{n})\mathbf{x}(p_{0}+p_{n})$ (2)
 Then we would like to capture the fine-grained spatial features as the source of key and value features.
 Since the central difference operations focus on the center-oriented gradients, they may overlook other texture patterns, which hinders the effectiveness of capturing manipulation traces.
-To obtain a diverse range of gradient information, we leverage the concept of Extended Local Binary Pattern (ELBP), which can encode the pixel relations from various perspectives, including central, angular, and radial directions Su et al.
-(2021).
+To obtain a diverse range of gradient information, we leverage the concept of Extended Local Binary Pattern (ELBP), which can encode the pixel relations from various perspectives, including central, angular, and radial directions [25].
 The angular direction calculates the difference between adjacent pixels in the outline of the kernel in a certain order, while the radial direction calculates the difference between radial adjacent pixels, see Fig.
-2.
+3.
 To obtain the key and value features, we adaptively ensemble these operations using a learnable coefficient vector, which is formulated as
-$\mathbf{x}^{K,V}(p_{0})=\sum_{p_{n}\in\mathcal{O}}\mathbf{w}(p_{n})\cdot$ $\left(\alpha_{c}\cdot\underbrace{(\mathbf{x}(p_{0}+p_{n})-\mathbf{x}(p_{0}))}_{\mathcal{D}_{c}}\right.$ (5)
-$+\alpha_{a}\cdot\underbrace{(\mathbf{x}(p_{0}+p_{n})-\mathbf{x}(p_{n-1}))}_{\mathcal{D}_{a}}$
-$+\alpha_{r}\cdot\left.\underbrace{(\mathbf{x}(p_{0}+p_{n})-\mathbf{x}(p_{\phi(n)}))}_{\mathcal{D}_{r}}\right)$
-$=\sum_{p_{n}\in\mathcal{O}}\mathbf{w}(p_{n})\cdot$ $\left(\underbrace{\left[\alpha_{c},\alpha_{a},\alpha_{r}\right]}_{\bm{\alpha}}\left[\begin{array}[]{l}\mathcal{D}_{c}\\ \mathcal{D}_{a}\\ \mathcal{D}_{r}\end{array}\right]\right).$
-In this equation, $\mathcal{D}_{c},\mathcal{D}_{a},\mathcal{D}_{r}$ denote central, angular, and radial difference operations respectively.
+$\mathbf{x}^{K,V}(p_{0})=\sum_{p_{n}\in\mathcal{O}}\mathbf{w}(p_{n})\left(\underbrace{\left[\alpha_{c},\alpha_{a},\alpha_{r}\right]}_{\bm{\alpha}}\left[\begin{array}[]{l}\mathcal{D}_{c}\\ \mathcal{D}_{a}\\ \mathcal{D}_{r}\end{array}\right]\right)$ (3)
+In this equation, $\mathcal{D}_{c},\mathcal{D}_{a},\mathcal{D}_{r}$ denote central, angular, and radial difference operations respectively, which can be defined as
+$\mathcal{D}_{c}$ $=\mathbf{x}(p_{0}+p_{n})-\mathbf{x}(p_{0}),$ (4)
+$\mathcal{D}_{a}$ $=\mathbf{x}(p_{0}+p_{n})-\mathbf{x}(p_{n-1}),$
+$\mathcal{D}_{r}$ $=\mathbf{x}(p_{0}+p_{n})-\mathbf{x}(p_{\phi(n)})$
 We use $3\times 3$ as the local receptive field $\mathcal{O}$ for instance and set $\mathcal{O}=\{(-1,-1),(-1,0),...,(1,1)\}$.
@@
 $\bm{\alpha}$ is the coefficient vector to balance the contribution of each operation.
-Given that different DeepFake face images may contain different manipulation traces, we design a learning scheme to adaptively balance the contribution strength of different DPDA operations.
-As our first branch concentrates on extracting coarse-grained information, its intermediate features can provide global instruction on the composition of manipulations.
-Denote the feature map of the corresponding block in our coarse-grained branch as $\mathbf{x}^{\prime}$.
-To achieve this adaptive learning, we design a sub-network $\mathcal{H}$ that contains several convolution layers.
-This network learns the $\bm{\alpha}$ based on $\mathbf{x}^{\prime}$ as
-$\bm{\alpha}={\rm Softmax}(\mathcal{H}(\mathbf{x}^{\prime})).$ (6)
+Adaptive Balance.
+Since various DeepFake faces exhibit diverse manipulation traces, the coefficient vector $\bm{\alpha}$ should be different.
+Thus, we design a strategy to dynamically regulate the influence of different DPDA operations.
+As the vanilla branch concentrates on global correlations, it has deeper insight into the manipulation composition.
+Hence, we set $\bm{\alpha}$ trainable and adapt it based on the knowledge from the vanilla branch.
+Let $\mathbf{x}^{\prime}$ be the intermediate feature of a block within the vanilla branch.
+We develop a sub-network $\mathcal{H}$ to learn $\bm{\alpha}$ as
+$\small\bm{\alpha}={\rm Softmax}(\mathcal{H}(\mathbf{x}^{\prime}))$ (5)
 After obtaining $\mathbf{x}^{Q}$ and $\mathbf{x}^{K,V}$, we project the query feature $\mathbf{x}^{Q}$ into query tokens, and project $\mathbf{x}^{K,V}$ into key and value tokens.
 Within each head, ${Q}_{i}=\mathbf{x}^{Q}{W}^{Q}_{i},{K}_{i}=\mathbf{x}^{K,V}{W}^{K}_{i},{V}_{i}=\mathbf{x}^{K,V}{W}^{V}_{i}$.
-Then the multi-head self-attention is calculated using the Eq.
-(2).
-3) Diversiform Pixel Difference Transformer Block.
-This block integrates the Diversiform Pixel Difference Attention (DPDA) module, which is similar to the Eq.
-(1).
-The overall process can be written as
-$\hat{\mathbf{x}}^{l}$ $=\mathbf{x}^{l}+{\rm DPDA}({\rm LN}(\mathbf{x}^{l})),$ (7)
-$\hat{\mathbf{x}}^{l+1}$ $=\hat{\mathbf{x}}^{l}+{\rm FFN}({\rm LN}(\hat{\mathbf{x}}^{l})),$
-With this strategy, this branch can effectively extract subtle manipulation traces, enhancing the performance in sequential detection.
-3.2 Sequential Decoder
-In order to predict sequential manipulations, we develop a sequential decoder architecture, which takes as input the features from the encoder and the embeddings of manipulation sequences.
-To effectively integrate these types of features into the decoder, we propose a Bidirectional Interactive Cross-attention (BICA) strategy that can properly integrate the features from the encoder into the decoder.
-Moreover, to improve the capacity of cross-attention, we introduce a shape-guided Gaussian mapping strategy to enhance its effectiveness.
-Spatial and Sequential Feature Fusing.
-To fully leverage the benefits of both coarse-grained and fine-grained features, we propose a Bidirectional Interaction Cross-attention (BICA) module.
-This module employs a hierarchical attention mechanism to explore the correlation between features and sequential annotations.
-1) Revisit of Conventional Cross-attention The cross-attention is typically employed in the decoder of the transformer for sequence tasks.
-Unlike self-attention mechanisms, the query, and key and value features usually originate from different sources.
-Given two input tokens from different sources $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$, the multi-head cross-attention (MCA) can be defined as
+Then the multi-head self-attention is calculated using
+$h_{i}={\rm Softmax}(Q_{i}K^{\top}_{i}/\sqrt{d_{i}})V_{i},$ (6)
+${\rm MSA}(\mathbf{x})={\rm Concat}(h_{1},...,h_{n})W^{o},$
+where $h_{i}$ is the self-attention of $i$-th head. Then we concatenate all these heads and perform a linear transformation with $W^{o}$ to obtain the final output.
+3.2 Multi-source Cross-attention (MSCA)
+Revisit of Conventional Cross-attention.
+Typically, the cross-attention [26] is employed for modeling such correlation [23].
+However, these mechanisms usually process the features from only two sources $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ (e.g., the encoder features and sequential embedding).
+The conventional Multi-head Cross-attention (MCA) can be defined as
 $Q_{i}=\mathbf{x}_{1}W^{Q}_{i},K_{i}=\mathbf{x}_{2}W^{K}_{i},V_{i}=\mathbf{x}_{2}W^{V}_{i},$ (7)
 $h_{i}={\rm Softmax}(Q_{i}K^{\top}_{i}/\sqrt{d_{i}})V_{i},$
-${\rm MCA}(\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{2})={\rm Concat}(h_{1},...,h_{n})W^{o},$
-2) Bidirectional Interactive Cross-attention.
+${\rm MCA}(\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{2})={\rm Concat}(h_{1},...,h_{n})W^{o}$
+Operations of MSCA.
 Let $\mathbf{x}^{c}$ and $\mathbf{x}^{f}$ represent the features from the coarse-grained and fine-grained branches.
-Let $\mathbf{x}^{s}$ be the embedding of sequential manipulation annotations.
+Let $\mathbf{x}^{s}$ be the embedding of inverted sequential manipulation annotations.
 In our task, it is intuitive to perform separate MCA operations on $\mathbf{x}^{s}$ with $\mathbf{x}^{c}$ and $\mathbf{x}^{f}$.
 However, this integration fails to explore deeper connections between these three features.
-Inspired by the concept of multi-modal cross-attentions Yu et al.
-(2019); Wei et al.
-(2020); Zhang et al.
-(2023), we propose Bidirectional Interactive Cross-attention (BICA) which employs a multi-level fusion strategy to fully excavate the underlying correlations.
-The overview of this module is shown in Fig.
-3.
-Figure 3: Overview of the sequential decoder (top). Illustration of Bidirectional Interactive Cross-attention (BICA). MCA denotes Multi-head Cross-attention.
-Firstly, we perform MSA on $\mathbf{x}^{s}$ to find the correlations of sequential manipulation annotations, which is then subjected to MCA with $\mathbf{x}^{c}$ and $\mathbf{x}^{f}$ to build the cross-attention.
-Then these processed features are summed up and applied LayerNorm to generate multi-modal features $\mathbf{a}^{c}$ and $\mathbf{a}^{f}$.
-These operations can be written as
-$\tilde{\mathbf{x}}^{s}$ $={\rm MSA}(\mathbf{x}^{s}),$ (9)
-$\tilde{\mathbf{a}}^{c}$ $={\rm MCA}(\tilde{\mathbf{x}}^{s},\mathbf{x}^{c},\mathbf{x}^{c}),\mathbf{a}^{c}={\rm LN}(\tilde{\mathbf{a}}^{c}+\tilde{\mathbf{x}}^{s}),$
-$\tilde{\mathbf{a}}^{f}$ $={\rm MCA}(\tilde{\mathbf{x}}^{s},\mathbf{x}^{f},\mathbf{x}^{f}),\mathbf{a}^{f}={\rm LN}(\tilde{\mathbf{a}}^{f}+\tilde{\mathbf{x}}^{s}).$
-We then perform interactive cross-attention between features of $\mathbf{a}^{c}$ and $\mathbf{a}^{f}$ in two directions as
-$\tilde{\mathbf{a}}^{c,f}$ $={\rm MCA}(\mathbf{a}^{c},\mathbf{a}^{f},\mathbf{a}^{f}),\mathbf{a}^{c,f}={\rm LN}(\tilde{\mathbf{a}}^{c,f}+\mathbf{a}^{c}),$ (10)
-$\tilde{\mathbf{a}}^{f,c}$ $={\rm MCA}(\mathbf{a}^{f},\mathbf{a}^{c},\mathbf{a}^{c}),\mathbf{a}^{f,c}={\rm LN}(\tilde{\mathbf{a}}^{f,c}+\mathbf{a}^{f}).$
-Next, we fuse the sequential features $\tilde{\mathbf{x}}^{s}$ with the enhanced spatial features $\mathbf{a}^{c,f}$ and $\mathbf{a}^{c,f}$, following the operations as
-$\mathbf{A}$ $=[\mathbf{a}^{c},\mathbf{a}^{f},\mathbf{a}^{c,f},\mathbf{a}^{f,c}],$ (11)
-$\tilde{\mathbf{x}}^{o}$ $={\rm MCA}(\tilde{\mathbf{x}}^{s},\mathbf{A},\mathbf{A}),$
-${\mathbf{x}}^{o}$ $={\rm LN}(\tilde{\mathbf{x}}^{o}+\tilde{\mathbf{x}}^{s}).$
-Through this fusion strategy, the sequential and spatial features can be fully interacted with and expose the underlying correlations.
-Shape-guided Gaussian Mapping In comparison to general image-sequence tasks, sequential manipulation annotations in our case are relatively short and may not provide sufficient guidance.
-To address this limitation, we propose a Shape-guided Gaussian Mapping strategy to leverage the strong spatial position priors associated with each manipulation component in the faces.
-Specifically, we assume that the position of each facial manipulation follows a Gaussian distribution.
-By learning these distributions, we can determine the probability of each position on a face image being manipulated.
-To simulate these Gaussian distributions, we construct a Variational Auto-encoder (VAE) Kingma and Welling (2013) with the image $\mathbf{I}$ as input.
-The VAE architecture contains an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$.
-The encoder transforms the input feature into means $\bm{\mu}$ and standards $\bm{\sigma}$ of these Gaussian distributions, while the decoder reconstructs the input images from the sampled latent code.
-Through the reconstruction process, the encoder learns to capture the spatial manipulation information and formulate them as Gaussian distributions.
-Next, we create a grid map $M$ that is composed of the $x$- and $y$- coordinates of each position.
-This grid map is then used to predict the probability of each position in different facial manipulations.
-These operations can be formulated as
-$\bm{\mu},\bm{\sigma}=\mathcal{E}(\mathbf{I}),$ (12)
-$S=\mathcal{N}(M\;|\;\bm{\mu},\bm{\sigma}),$
-where $S$ is the probability map for each spatial position.
-This map is integrated into the MCA operations ${\rm MCA}(\tilde{\mathbf{x}}^{s},\mathbf{x}^{c},\mathbf{x}^{c})$ and ${\rm MCA}(\tilde{\mathbf{x}}^{s},\mathbf{x}^{f},\mathbf{x}^{f})$ in Eq.
-(9) respectively.
-We take the integration in ${\rm MCA}(\tilde{\mathbf{x}}^{s},\mathbf{x}^{c},\mathbf{x}^{c})$ for example as
-$Q_{i}=\tilde{\mathbf{x}}^{s}W^{Q}_{i},K_{i}=\mathbf{x}^{c}W^{K}_{i},V_{i}=\mathbf{x}^{c}W^{V}_{i},$ (13)
-$h_{i}={\rm Softmax}(Q_{i}K^{\top}_{i}/\sqrt{d_{i}}+\eta\cdot S)V_{i},$
-${\rm MCA}(\tilde{\mathbf{x}}^{s},\mathbf{x}^{c},\mathbf{x}^{c})={\rm Concat}(h_{1},...,h_{n})W^{o}.$
-By inserting the shape priors of facial manipulation, the cross-attention is greatly enhanced to capture the spatial and sequential forgery traces.
-Figure 4: The impact of the sequential manipulation on facial attributes. The second manipulation impacts the first manipulation on (left) eye and (right) eyeglasses.
-Inverted Order Prediction.
-Based on our analysis, we invert the annotation of the manipulation sequence for training.
-For instance, given a manipulated face with annotation of “eyes-nose-eyebrow-lips-hair”, we reorder it as “hair-lips-eyebrow-nose-eyes”, and convert it to $\mathbf{x}^{s}$ for sequential modeling.
-By doing so, the model first predicts the last manipulation of “hair”.
-By isolating the disturbance caused by “hair”, the model can then focus on predicting the next manipulation, which is “lips”.
-The rationale behind this strategy is that the manipulation of the nose can affect the manipulation of the eyes, making the detection of eye manipulations more challenging.
-On the other hand, the manipulation in hair is not influenced by other manipulations, making it more easily detectable (see Fig.
-4).
-The advantage of this strategy is that, compared to forward sequence predictions, each step in the inverted sequence predictions is only influenced by itself and is independent of other types of manipulations.
-3.3 Loss Function
-In the training phase, we employ an auto-regressive mechanism Graves (2013) to model sequence relationships with the cross-entropy loss $L_{\rm ce}$.
-This loss measures the error of predicted manipulation sequence and ground truth.
-For the shape-guided module, we perform the general image reconstruction loss $L_{\rm rec}$ and KLD loss $L_{\rm kld}$ used in VAE Kingma and Welling (2013).
-The overall loss function is expressed as
-$L=L_{\rm ce}+\lambda_{1}L_{\rm rec}+\lambda_{2}L_{\rm kld},$ (14)
-where ${\lambda}_{1},{\lambda}_{2}$ are weights for balancing these loss terms.
+As such, we propose a Multi-source Cross-attention (MSCA) module to capture the correlations among multiple sources.
+The overview of the entire module is illustrated in Fig.
+4.
+Firstly, we concatenate coarse-grained features $\mathbf{x}^{c}$ and fine-grained features $\mathbf{x}^{f}$ along the channel dimension.
+Subsequently, we facilitate interactions between the two different levels of fabricated features throuth an MLP structure, yielding shallowly fused sequencetial features $\mathbf{x}^{m}$.
+Then, by performing self-attention computations on $\mathbf{x}^{s}$, we extract features that exploit correlations among annotations of sequence tampering.
+These operations can be formulated as follows:
+$\small\mathbf{x}^{m}={\rm MLP}({\rm Concat}(\mathbf{x}^{c},\mathbf{x}^{f})),\;\tilde{\mathbf{x}}^{s}={\rm MSA}(\mathbf{x}^{s})$ (8)
+Next, we implement a second-stage fusion strategy, performing multi-head cross-attention between $\tilde{\mathbf{x}}^{s}$ and $\mathbf{x}^{c}$, $\mathbf{x}^{m}$, $\mathbf{x}^{f}$ respectively, as:
+$\tilde{\mathbf{a}}^{c}$ $={\rm MCA}(\tilde{\mathbf{x}}^{s},\mathbf{x}^{c},\mathbf{x}^{c}),\mathbf{a}^{c}={\rm LN}(\tilde{\mathbf{a}}^{c}+\tilde{\mathbf{x}}^{s}),$ (9)
+$\tilde{\mathbf{a}}^{m}$ $={\rm MCA}(\tilde{\mathbf{x}}^{s},\mathbf{x}^{m},\mathbf{x}^{m}),\mathbf{a}^{m}={\rm LN}(\tilde{\mathbf{a}}^{m}+\tilde{\mathbf{x}}^{s}),$
+$\tilde{\mathbf{a}}^{f}$ $={\rm MCA}(\tilde{\mathbf{x}}^{s},\mathbf{x}^{f},\mathbf{x}^{f}),\mathbf{a}^{f}={\rm LN}(\tilde{\mathbf{a}}^{f}+\tilde{\mathbf{x}}^{s})$
+Finally, we directly sum up the three types of features and then pass them through an MLP to filter out irrelevant noise, thereby uncovering identical manipulation patterns and obtaining the final sequential deepfake features, following the operations as
+$\small\mathbf{A}=\tilde{\mathbf{a}}^{c}+\tilde{\mathbf{a}}^{m}+\tilde{\mathbf{a}}^{f},\;\tilde{\mathbf{x}}^{o}={\rm MLP}(A),\;\mathbf{x}^{o}={\rm LN}(\tilde{\mathbf{x}}^{o}+\mathbf{A})$ (10)
+Figure 5: Integration with MSCA.
 Table 1: Performance of different methods on detecting sequential DeepFakes.
 Methods Facial-components Facial-attributes Average
-Fixed-Acc Adaptive-Acc Fixed-Acc Adaptive-Acc Fixed-Acc Adaptive-Acc
-Multi-Cls (ResNet-34) 69.66 50.54 66.99 46.68 68.33 48.61
-Multi-Cls (ResNet-50) 69.65 50.57 66.66 46.00 68.16 48.29
-DETR (ResNet-34) 69.87 50.63 67.93 48.15 68.90 49.39
-DETR (ResNet-50) 69.75 49.84 67.62 47.99 68.69 48.92
-DRN (ResNet-34) 66.06 45.79 64.42 43.20 65.24 44.50
-MA (ResNet-34) 71.31 52.94 67.58 47.48 66.45 50.21
-Two-Stream (ResNet-34) 71.92 53.89 66.77 46.38 69.35 50.14
-SeqFake-Former (ResNet-34) 70.88 52.96 67.85 48.36 69.37 50.66
-SeqFake-Former (ResNet-50) 69.41 50.22 68.38 48.97 68.90 49.60
-TS-Former (Ours) (ResNet-34) 74.70 58.43 69.13 49.68 71.92 54.06
-TS-Former (Ours) (ResNet-50) 73.32 56.08 68.86 49.67 71.09 52.88
-Table 2: Ablation study on the facial-components datasets.
-Setting Resnet-34 Resnet-50
-Fixed-Acc Adaptive-Acc Fixed-Acc Adaptive-ACC
-Baseline 69.65 (+0.00) 50.43 (+0.00) 68.84 (+0.00) 49.40 (+0.00)
-Baseline+DPDA 70.24 (+0.59) 51.87 (+1.44) 69.23 (+0.39) 51.05 (+1.65)
-Baseline+DPDA+BICA 71.00 (+1.35) 52.67 (+2.24) 71.55 (+2.71) 53.06 (+3.66)
-Baseline+DPDA+BICA+IOP 73.32 (+3.67) 56.65 (+6.22) 72.65 (+3.81) 55.03 (+5.63)
-Baseline+DPDA+BICA+IOP+SGM 74.70 (+5.05) 58.43 (+8.00) 73.32 (+4.48) 56.08 (+6.68)
+FACC AACC FACC AACC FACC AACC
+Multi-Cls (ResNet-34) [28] 69.66 50.54 66.99 46.68 68.33 48.61
+Multi-Cls (ResNet-50) [28] 69.65 50.57 66.66 46.00 68.16 48.29
+DETR (ResNet-34) [2] 69.87 50.63 67.93 48.15 68.90 49.39
+DETR (ResNet-50) [2] 69.75 49.84 67.62 47.99 68.69 48.92
+DRN [29] 66.06 45.79 64.42 43.20 65.24 44.50
+DRN [29] 69.38 49.76 67.66 47.31 68.52 48.54
+MA [39] 71.31 52.94 67.58 47.48 69.45 50.21
+MA [39] 70.12 51.82 68.91 49.19 69.52 50.51
+Two-Stream [18] 71.92 53.89 66.77 46.38 69.35 50.14
+Two-Stream [18] 72.00 53.97 65.20 44.39 68.60 49.18
+SeqFake-Former (ResNet-34) [23] 72.13 54.80 67.99 48.32 70.06 51.56
+SeqFake-Former (ResNet-34) [23] 70.97 52.65 68.40 48.93 69.69 50.79
+SeqFake-Former (ResNet-50) [23] 72.65 55.30 68.86 49.63 70.76 52.47
+SeqFake-Former (ResNet-50) [23] 69.12 50.03 67.93 48.65 68.53 49.34
+MMNet (ResNet-50) [31] 73.93 56.83 69.27 50.44 71.60 53.64
+TSOM (ResNet-34) 75.21 59.74 69.18 50.54 72.20 55.14
+TSOM (ResNet-50) 75.53 59.67 69.80 51.21 72.67 55.44
+Table 2: Performance of different methods on challenging scenarios.
+Methods Facial-components Facial-attributes Average
+FACC AACC FACC AACC FACC AACC
+Two-Stream 24.00 12.74 31.55 16.92 27.78 14.83
+DRN 48.11 14.11 50.35 20.62 49.23 17.37
+MA 59.08 33.38 67.20 47.07 63.14 40.23
+SeqFake-Former (ResNet-34) 58.70 31.43 62.88 41.88 60.79 36.66
+SeqFake-Former (ResNet-50) 57.35 29.50 64.81 44.49 61.08 37.00
+TSOM (ResNet-34) 61.79 36.80 67.44 47.14 64.62 41.97
+TSOM (ResNet-50) 62.52 37.21 67.33 48.03 64.93 42.62
+3.3 Shape-guided Gaussian Mapping
+We propose a Shape-guided Gaussian Mapping (SGM) strategy to leverage the strong spatial position priors associated with each manipulation component in the faces.
+Since each manipulation has its specific characteristic, its manipulation shape likely follows a certain distribution.
+By learning these distributions, we can predict the probability of each position on a face image being manipulated.
+This probability can serve as shape priors to further enhance the effectiveness of MSCA.
+Distribution Estimation.
+Since no ground truth is provided, estimating these distributions is non-trivial.
+Inspired by unsupervised learning strategies, we construct a Variational Auto-encoder (VAE) [11] to estimate these distributions.
+Specifically, we construct an encoder $\mathcal{E}$, which transforms the input image $\bm{\mathcal{I}}$ into means $\bm{\mu}$ and standards $\bm{\sigma}$ of these distributions as $\bm{\mu},\bm{\sigma}=\mathcal{E}(\bm{\mathcal{I}})$.
+Note that $\bm{\mu},\bm{\sigma}$ are vectors containing mean and standard of each manipulation.
+To ensure these distributions align with the input image, we employ a decoder $\mathcal{D}$ for reconstruction based only on the estimated distributions.
+Integration with MSCA.
+We create a grid map $M$ that is composed of the $x$- and $y$- coordinates of each position.
+This grid map is then used to predict the probability of each position in different facial manipulations.
+This process can be formulated as $S=\mathcal{N}(M\;|\;\bm{\mu},\bm{\sigma})$, where $S$ is the probability map for each spatial position.
+Considering different heads in MSCA have different focuses.
+To properly integrate SGM with MSCA, we craft different probability maps $\{S^{\prime}_{i}\}_{i=1}^{h}$ for each head.
+Specifically, we use $\bm{\mu},\bm{\sigma}$ as the base of all heads and predict the head-specific offset $\Delta\bm{\mu}_{i},\Delta\bm{\sigma}_{i}$ for $i$-th head.
+These offsets are obtained as $\Delta\bm{\mu}_{i},\Delta\bm{\sigma}_{i}={\rm FC}(\mathcal{E}_{f}(\bm{\mathcal{I}}))$, where $\mathcal{E}_{f}(\bm{\mathcal{I}})$ denotes the last feature of encoder $\mathcal{E}$.
+Therefore, the probability map $S^{\prime}_{i}$ for $i$-th head can be defined as
+$\small S^{\prime}_{i}=\mathcal{N}(M\;|\;\bm{\mu}+\Delta\bm{\mu_{i}},\bm{\sigma}+\Delta\bm{\sigma_{i}})$ (11)
+This process is illustrated in Fig.
+5.
+Then these maps are integrated into ${\rm MCA}(\tilde{\mathbf{x}}^{s},\mathbf{x}^{c},\mathbf{x}^{c})$ and ${\rm MCA}(\tilde{\mathbf{x}}^{s},\mathbf{x}^{f},\mathbf{x}^{f})$ operations in MSCA, by calculating the attention as
+$\small h_{i}={\rm Softmax}(Q_{i}K^{\top}_{i}/\sqrt{d_{i}}+\eta S^{\prime}_{i})V_{i}$ (12)
+3.4 Inverted Order Prediction
+In the training phase, we invert the annotation of the manipulation sequence.
+For instance, as shown in Fig.
+4, the manipulated face is annotated as “eye-nose-eyebrow-hair-lip”, we rearrange it to ”lip-hair-eyebrow-nose-eye” as the input of the MSCA module for sequential modeling.
+This strategy enables the model to first predict the final manipulation (e.g., hair) and understand the consequent disturbance caused by it.
+Then the model concentrates on the preceding manipulation (e.g., lips).
+The rationale behind this strategy is that the subsequent manipulation (e.g., noise) may affect the preceding ones (e.g., eye), making it more challenging to detect them (recall Fig.
+2).
+In contrast, preceding manipulation (e.g., hair) is not influenced by subsequent ones, making it more easily detectable.
+In summary, this strategy offers the advantage that, compared to forward sequence predictions, each step in the inverted sequence predictions is only influenced by itself and remains independent of other manipulations.
+3.5 Objectives
+In the training phase, the auto-regressive mechanism [5] is employed to model sequence relationships with the cross-entropy loss $L_{\rm ce}$.
+This loss measures the error of the predicted manipulation sequence and ground truth.
+To learn the shape priors, we perform the image reconstruction loss $L_{\rm rec}$ and KLD loss $L_{\rm kld}$ on constructed auto-encoder $(\mathcal{E},\mathcal{D})$ as in [11].
+The overall loss function is expressed as
+$L=L_{\rm ce}+\lambda_{1}L_{\rm rec}+\lambda_{2}L_{\rm kld}.$ (13)
+where ${\lambda}_{1},{\lambda}_{2}$ are weights for balancing these loss terms.
+Table 3: Effect of each component.
+Setting Resnet-34 Resnet-50 Average
+FACC AACC FACC AACC FACC AACC
+BL 70.45 (+0.00) 52.18 (+0.00) 69.47 (+0.00) 50.89 (+0.00) 69.96 (+0.00) 51.34 (+0.00)
+BL+DPDA 71.07 (+0.62) 52.84 (+0.66) 71.06 (+1.59) 52.80 (+1.91) 71.07 (+1.11) 52.82 (+1.48)
+BL+DPDA+MSCA 71.74 (+1.29) 54.05 (+1.87) 72.04 (+2.57) 54.42 (+3.53) 71.89 (+1.93) 54.24 (+2.90)
+BL+DPDA+MSCA+IOP 74.22 (+3.77) 58.19 (+6.01) 74.30 (+4.83) 58.05 (+7.16) 74.26 (+4.30) 58.12 (+6.78)
+BL+DPDA+MSCA+IOP+SGM 75.21 (+4.76) 59.74 (+7.56) 75.53 (+6.06) 59.67 (+8.78) 75.37 (+5.41) 59.71 (+8.37)
+Table 4: Performance of applying IOP on various methods.
+Methods Facial-components Facial-attributes Average
+FACC AACC FACC AACC FACC AACC
+Two-Stream 72.00 53.97 65.20 44.39 68.60 49.18
+Two-Sream+IOP 73.55(+1.55) 58.03(+4.06) 67.60(+2.40) 47.78(+3.39) 70.58(+1.98) 52.91(+3.73)
+DRN 69.38 49.76 67.66 47.31 68.52 48.54
+DRN+IOP 71.81(+2.43) 54.21(+4.45) 68.89(+1.23) 49.74(+2.43) 70.35(+1.83) 51.98(+3.44)
+MA 70.12 51.82 68.91 49.19 69.52 50.51
+MA+IOP 73.41(+3.29) 57.30(+5.48) 68.13(-0.78) 48.58(-0.61) 70.77(+1.25) 52.94(+2.43)
+SeqFake-Former 70.97 52.65 68.40 48.93 69.69 50.79
+SeqFake-Former+IOP 73.34 (+2.37) 56.93 (+4.28) 69.09(+0.69) 48.40 (-0.53) 71.22 (+1.53) 52.67(+1.88)
 4 Experiments
@@
 Datasets.
-Our experiments are conducted on Sequential DeepFake Dataset (SDD) Shao et al.
-(2022).
+To the best of our knowledge, there is only one public dataset for sequential DeepFake detection that proposed in [23].
 This dataset contains two types of sequential DeepFake manipulations, the facial components manipulation and the facial attributes manipulation.
@@
 Evaluation Metrics.
-In line with previous work Shao et al.
-(2022), we employ Fixed Accuracy (Fixed-Acc) and Adaptive Accuracy (Adaptive-Acc) for evaluation.
-To calculate the Fixed-Acc score, the sequence annotations shorter than five are padded with the label “no manipulation” and then compare the predictions with the padded sequence annotation.
+In line with previous work [23], we employ Fixed Accuracy (FACC) and Adaptive Accuracy (AACC) for evaluation.
+To calculate the FACC score, the sequence annotations shorter than five are padded with the label “no manipulation” and then compare the predictions with the padded sequence annotation.
 On the other hand, due to the nature of the auto-regress mechanism, the model can stop predicting upon encountering the End-of-Sequence (EOS) symbol, resulting in various lengths of the predicted sequence.
-The adaptive-Acc score is calculated by comparing the predictions with the sequence annotations in terms of dynamic lengths.
+The AACC score is calculated by comparing the predictions with the sequence annotations in terms of dynamic lengths.
 Implementation Details.
-We utilize ResNet-34 He et al.
-(2015) as our CNN stem in main experiment.
 Each branch in the spatial encoder has two Transformer blocks.
@@
 4.2 Results
-Table 1 shows the performance of our method compared to other counterparts, including Multi-Cls, DETR Carion et al.
-(2020), DRN Wang et al.
-(2019), MA Zhao et al.
-(2021b), Two-Stream Luo et al.
-(2021), SeqFake-Former Shao et al.
-(2022) respectively.
+Table 1 compares the performance of our method with other counterparts, including Multi-Cls, DETR [2], DRN [29], MA [39], Two-Stream [18], SeqFake-Former [23] respectively.
 Multi-Cls and DETR serve as baseline methods adapted from models for general vision tasks.
@@
 DRN, MA, and Two-Stream are DeepFake detection methods adapted by substituting the binary classification head with a multi-classification head.
-These methods are all referrenced from Shao et al.
-(2022).
-SeqFake-Former is a recent dedicated method for sequential DeepFake detection.
-We reproduce this method rigorously using their release codes.
-As shown in Table 1, our method outperforms baseline methods by a large margin, averaging $71.92\%$ and $54.06\%$ in Fixed-Acc and Adaptive-Acc scores, respectively.
-Moreover, our method surpasses the DeepFake detection methods in both tracks.
-Notably, compared to SeqFakeFormer, our method achieves improvements by $3.82\%,5.47\%$ and $1.28\%,1.32\%$ in these two tracks.
-The results also indicate that all methods perform better in Fixed-Acc than in Adaptive-Acc.
-This observation is consistent with the findings in Shao et al.
-(2022) that predicting manipulation sequences with varying lengths is a more challenging task.
-In real-world applications, manipulation sequences with varying lengths are more practical, thus the improvement of our method in this metric sufficiently represents the practicability of the proposed modules.
-Intriguingly, we observe that ResNet-50 performs slightly worse than ResNet-34 across all methods.
-This contrasts with the expectation that a larger model should have stronger capacity.
-We discuss this phenomenon in Sec.
-4.4.
-4.3 Ablation Study
-Our method consists of several key components, including the Diversified Pixel Difference Attention module (DPDA), Bidirectional Interactive Cross-attention (BICA), Inverted Order Prediction (IOP), and Shape-guided Gaussian Mapping (SGM).
+SeqFake-Former [23] is a dedicated method for sequential DeepFake detection.
+MMNet [31] is the latest work achieving state-of-the-art performance.
+The results of Multi-Cls, DETR, DRN, MA, and Two-stream are referred from [23].
+Note that only DRN, MA, Two-stream, and SeqFake-Former have released their codes.
+For better evaluation, we reproduce them with their released codes (marked by *) (More details in Supplementary).
+As shown in Table 1, our method outperforms others in both tracks, averaging $72.67\%$ and $55.44\%$ in FACC and AACC scores on ResNet-50.
+Compared to SeqFake-Former, our method improves by $2.88\%,4.37\%$ and $0.94\%,1.58\%$ in these two tracks on ResNet-50.
+Moreover, our method averagely surpasses MMNet by $1.07\%$ and $1.8\%$ in FACC and AACC.
+In Challenging Scenarios.
+In real-world scenarios, images often undergo various post-processing operations.
+To these scenarios, we perform post-processing operations on the facial-components and facial-attributes datasets.
+Specifically, for each image, we randomly select one or two processes from a set that includes adding Gaussian noises, shifting RGB channels, compression, converting to gray, and jittering colors.
+This increases the difficulty of original datasets, which we refer to as Facial-components and Facial-attributes.
+We then evaluate different methods under these challenging scenarios.
+As shown in Table 2, while all methods suffer a notable performance drop, our model still outperforms others, demonstrating its robustness against external disturbances.
+4.3 Analysis
+Effect of Each Component.
+Our method consists of several key components, including the Diversified Pixel Difference Attention module (DPDA), Multi-source Cross-attention (MSCA), Inverted Order Prediction (IOP), and Shape-guided Gaussian Mapping (SGM).
 To evaluate the effectiveness of each component, we conduct a series of ablation studies on the facial-components track using the stem of ResNet-34 and ResNet-50.
-Specifically, we designed different models based on the following settings: 1) A baseline model with a Transformer architecture that only consists of basic encoders and decoders.
-2) Adding the DPDA module to the baseline model, which generates two spatial features from the encoder.
-Then we directly add these two features and feed them into the sequential decoder.
-3) Adopting the BICA strategy on the baseline model for feature integration.
+Specifically, 1) the baseline model (BL) contains only basic encoders and decoders.
+2) Adding the DPDA module to the baseline model and directly feeding the features from the encoder into the sequential decoder.
+3) Employ the MSCA strategy for feature integration.
 4) Incorporating inverted order prediction.
-5) Inserting the SGM module into BICA strategy.
-The results of the ablation studies are shown in Table 2.
-It can be observed that each component contributes to the overall performance.
-Taking Resnet-34 as an example.
-The DPDA module improves the Fixed-Acc and Adaptive-Acc by $0.59\%$ and $1.44\%$, respectively.
-The BICA strategy further enhances the performance by $0.76\%$ and $0.8\%$ in Fixed-Acc and Adaptive-Acc.
-After applying IOP, the performance is significantly improved by $2.32\%$ in Fixed-Acc and $3.98\%$ in Adaptive-Acc, showing the effectiveness of this strategy in gradually capturing manipulation traces.
-Finally, the integration of SGM yields further performance, achieving the best performance when all components are applied.
-Figure 5: The visualization of our model and SeqFakeFormer under the (a) eyebrow-eye-hair-nose-lip and (b) nose-eyebrow-lip-eye-hair manipulating annotations.
-Table 3: Performance on Different CNN Stems.
-Stem Fixed-Acc Adaptive-Acc
-SFF Ours SFF Ours
-ResNet-18 68.41 69.97 49.55 52.26
-ResNet-34 70.88 74.70 52.96 58.43
-Resnet-50 69.41 73.32 50.22 56.08
-Resnet-101 68.87 73.46 49.06 56.77
-4.4 Discussion
+5) Inserting the SGM module into the MSCA strategy.
+The results shown in Table 3 indicate the contribution of each component to the overall performance.
+For instance, with Resnet-34, the DPDA module improves the FACC and AACC by $0.59\%$ and $1.44\%$, respectively.
+The MSCA strategy further enhances the performance by $0.76\%$ in FACC and $0.8\%$ in AACC.
+The IOP significantly boosts FACC by $2.32\%$ and AACC by $3.98\%$, showing its effectiveness in gradually capturing manipulation traces.
+Finally, the integration of all components leads to the best performance.
+Generalizability of IOP.
+The proposed IOP strategy is applicable to many architectures.
+To verify this, we evaluate the performance of different methods with/without IOP.
+As shown in Table 4, all methods have gained enhancement on average.
+Even the methods of MA, DRN, and two-stream that are not specifically designed for sequence detection tasks surprisingly improve their performance by a notable margin.
+In particular, by applying IOP, Two-Stream even achieves competitive performance with MMnet on the facial-components track, demonstrating the generic effect of IOP for SDD.
+Table 5: Different loss weights $\lambda_{1},\lambda_{2}$.
+$\lambda_{1}$ $\lambda_{2}$ FACC AACC
+0.1 0.1 72.21 56.02
+0.01 0.01 73.32 56.84
+0.5 0.01 74.20 58.21
+1 0.01 73.71 57.24
+0.1 0.01 75.21 59.74
+Study on Loss Weights $\lambda_{1},\lambda_{2}$.
+We study the effect of various loss weight $\lambda_{1}$ and $\lambda_{2}$.
+The results are shown in Table 5.
+It can be seen that our method is not very sensitive to the values of loss weights (±$(\pm 3\%)$.
+This is because the classification loss primarily influences the learning direction during training.
+Compared to the KLD loss $\lambda_{2}$, the reconstruction loss $\lambda_{1}$ seemly has more impact, as it directly instructs the training of encoder ${\cal E}$ in a self-supervised manner.
+In the main experiment, we use $\lambda_{1}=0.1,\lambda_{2}=0.01$.
+Adaptation on One-step DeepFake Detection (ODD).
+Since one-step detection does not require sequential clues, existing SDD methods can not be directly applied to this task.
+For more comprehensive studies, we adapt these methods by replicating the one-step label $N=5$ times as the sequential annotations, where $N$ is the maximum manipulation length in dataset.
+For example, the real faces are assigned with annotations of $[0,0,0,0,0]$, while forged faces correspond to $[1,1,1,1,1]$.
+For comparison, we train these methods on the FF++ dataset [21] under this configuration.
+Note the basenetwork of SeqFake-Former and ours is ResNet-34.
+The results are shown in Table 6, revealing that our method still achieves the best results.
+Table 6: Performance on FF++.
+Methods ACC
+Two-stream 82.40
+DRN 81.42
+MA 82.64
+SeqFake-Former 82.43
+TSOM 83.70
+Figure 6: Attention visualization.
 Attention Visualization.
-To assess the learning ability of our method, we visualize the attention maps of our method compared with Seq-FakeFormer under two different manipulation annotations, eyebrow-eye-hair-nose-lip and nose-eyebrow-lip-eye-hair.
 Fig.
-5 shows the Grad-CAM Selvaraju et al.
-(2019) of these methods.
+6 shows the Grad-CAM [22] of different methods.
 We observe that under the same annotation, our method can identify the manipulated regions more precisely.
@@
 Our method can capture this subtle alternation as evidence for exposing the sequential manipulations.
-Controversy Regarding CNN Stem.
-From Table 3 we observe an unexpected phenomenon where using ResNet-34 as CNN stem achieves better performance than ResNet-50.
-To further investigate this, we validate our method using ResNet-18 and ResNet-101.
-Table 3 shows the performance comparison of our method and SeqFake-Former (SFF) on these architectures.
-It can be seen that the performance of ResNet-18 is notably degraded compared to ResNet-34, aligning with our expectations.
-However, when using ResNet-101 as the Stem, it yields lower performance than ResNet-50.
-This finding contradicts the general understanding that increasing the model size and parameters can improve feature extraction capacity.
-One possible explanation for this phenomenon is that the length of the manipulation sequence is relatively short (less than five), which increases the likelihood of overfitting in large models.
-This suggests the need to strike a balance between model size and performance in the future.
 5 Conclusion
-We introduce a novel Texture-aware and Shape-guided Transformer to enhance detection performance.
-Compared to the classic Transformer-based methods, we introduce four improvements.
-Firstly, we describe a new Diversiform Pixel Difference Attention module (DPDA) to capture subtle manipulation traces.
-Moreover, we introduce a Bidirectional Interaction Cross-attention module (BICA) to explore deep correlations between spatial and sequential features.
-To further enhance the cross-attention, we describe a Shape-guided Gaussian mapping strategy to provide the priors of manipulation shapes.
-Lastly, we improve the performance by inverting the prediction order from forward to backward.
-The results obtained on public datasets demonstrate the efficacy of our method in sequential DeepFake detection.
+This paper introduces a new Transformer design, called TSOM, for sequential DeepFake detection.
+Our method is inspired by three perspectives: Texture, Shape, and Order of Manipulations, leading to four key improvements.
+Firstly, we propose a text-aware branch featuring new Diversiform Pixel Difference Attention modules (DPDA) to capture subtle manipulation traces.
+Secondly, we introduce a Multi-source Cross-attention module (MSCA) to explore deep correlations between spatial and sequential features.
+To enhance the cross-attention further, we describe a Shape-guided Gaussian Mapping (SGM) strategy to incorporate priors on manipulation shapes.
+Lastly, we improve performance using a simple Inverted Order prediction (IOP) strategy, which reverses the manipulation annotation order from forward to backward.
+The results on public datasets demonstrate the efficacy of our method in sequential DeepFake detection.
+Acknowledgement.
+This work is supported in part by the National Natural Science Foundation of China (No.62402464), Shandong Natural Science Foundation (No.ZR2024QF035), China Postdoctoral Science Foundation (No.2021TQ0314, No.2021M703036).